### v1

# Primes and The Field of Values of Characters

Abstract.
Let $p$ be a prime. For $p=2$, the fields of values of the complex irreducible characters of finite groups whose degrees are not divisible by $p$ have been classified; for odd primes $p$, a conjectural classification has been proposed. In this work, we extend this conjecture to characters whose degrees are divisible by arbitrary powers of $p$, and we provide some evidence supporting its validity.
Key words and phrases:
Character, field of values, conductor
2020 Mathematics Subject Classification:
Primary 20C15; Secondary 20C30, 20C33
The first author gratefully acknowledges the support of the AMS-Simons Research Enhancement Grant (AWD-000167 AMS). The work of the second author is supported by Grant PID2022-137612NB-I00 funded by MCIN/AEI/ 10.13039/501100011033 and ERDF “A way of making Europe.” The third author gratefully acknowledges the support of the NSF (grant DMS-2200850) and the Joshua Barlaz Chair in Mathematics.
1. Introduction
In [ILNT19], Isaacs et al. discovered an unexpected connection between the irrationalities of an irreducible character of a finite group and the parity of its degree: if $\chi$ is an irreducible character of odd degree and its field of values is ${\mathbb{Q}}(\sqrt{d})$ for some square-free integer $d\neq-1$, then $d\equiv 1\pmod{4}$. (Here, the field of values of $\chi$, denoted ${\mathbb{Q}}(\chi)$, is the smallest extension of $\mathbb{Q}$} containing all values of $\chi$.) This result was later generalized to arbitrary higher irrationalities by the second and third authors in [NT21], providing a complete classification of the fields of values of the irreducible characters of odd degree. For odd primes $p$, this description is conjectured, remains open, and is reduced in [NT21] to a problem on quasi-simple groups.
Needless to say, the focus so far has been on characters of degree not divisible by $p$ (or on height zero characters [NRTV24], if one introduces Brauer $p$-blocks), largely motivated by the Galois–McKay conjecture ([N04]) and its possible extensions. An apparently innocent question raised in [IN24] (Question D), however, has shifted this perspective: could the results of [NT21] also be extended to characters of degree divisible by $p$? Given a prime $p$, what are the field of values of the characters divisible by a given power of $p$? The most natural setting to look for examples and patterns is when $p=2$ in quadratic fields ${\mathbb{Q}}(\sqrt{d})$, where $d$ is a square-free integer. In all the cases that we have checked, we observe, for instance, that if ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\sqrt{d})$ with even $d\neq\pm 2$, then $\chi(1)$ is divisible by $4$.
Our first result solves one direction of the classification of the abelian number fields that can occur as fields of values of irreducible characters whose degree has a given $p$-part. Throughout this paper, let $\nu_{p}$ denote the $p$-adic valuation function on the integers; that is, $\nu_{p}(n)=a$ if $p^{a}$ is the largest power of $p$ dividing $n$. If $\mathbb{F}/{\mathbb{Q}}$ is an abelian extension, recall that the conductor $c(\mathbb{F})$ is the smallest integer $n\geq 1$ such that $\mathbb{F}$ is contained in the $n$-th cyclotomic field ${\mathbb{Q}}_{n}:={\mathbb{Q}}(e^{2\pi i/n})$. Also, if $m$ is any positive integer, we let $\mathbb{F}_{m}:=\mathbb{F}(e^{2\pi i/m})$. For a finite group $G$ and $\chi\in{\rm Irr}(G)$, we define $c(\chi)$ to be the conductor of the field ${\mathbb{Q}}(\chi)$.
THEOREM A.
Let $p$ be a prime, $\mathbb{F}$ an abelian extension of $\mathbb{Q}$}, and $a:=\nu_{p}(c(\mathbb{F}))$. Then, for every integer $b\geq\nu_{p}([\mathbb{F}_{p^{a}}:\mathbb{F}])$, there exist a finite group $G$ and $\chi\in\operatorname{Irr}(G)$ such that ${\mathbb{Q}}(\chi)=\mathbb{F}$ and $\nu_{p}(\chi(1))=b$.
To complete the classification, we would need to prove the following, which extends the still unproved Conjecture B.3 of [NT21] on characters of degree not divisible by $p$ to arbitrary characters.
CONJECTURE B.
Let $\chi$ be an irreducible character of some finite group such that $a:=\nu_{p}(c(\chi))\geq 1$. Then $\chi(1)$ is divisible by ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$.
A few remarks are in order. First, the degree of the extension ${\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi)$ can be any power of $p$ when $a$ varies. Indeed: on one hand, every abelian extension $\mathbb{F}/{\mathbb{Q}}$ can arise as a character field (see Lemma 2.3). On the other hand, given a prime power $p^{a}$, elementary number theory shows that there exists an extension $\mathbb{F}\supseteq{\mathbb{Q}}_{p}$ such that $\nu_{p}(c(\mathbb{F}))=a$ and $\mathbb{F}_{p^{a}}:\mathbb{F}]=p^{a-1}$. Second, if ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\sqrt{d})$ with even $d\neq\pm 2$, then $c(\chi)=4|d|$. Hence $\nu_{2}(c(\chi))=3$ and ${\mathbb{Q}}_{2^{3}}(\chi):{\mathbb{Q}}(\chi)]=4$. This would explain our earlier observation on the quadratic character fields and the divisibility by $4$ of character degrees.
Since ${\mathbb{Q}}_{p}(\chi):{\mathbb{Q}}(\chi)]=[{\mathbb{Q}}_{p}:({\mathbb{Q}}_{p}\cap{\mathbb{Q}}(\chi))]$, which divides ${\mathbb{Q}}_{p}:{\mathbb{Q}}]=p-1$, it follows that ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$ is precisely the $p$-part of ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}(\chi)]$. Theorem A therefore implies the following.
THEOREM C.
Let $\mathbb{F}$ be an abelian extension of $\mathbb{Q}$} and $a:=\nu_{p}(c(\mathbb{F}))$. Let $b\in{\mathbb{Z}}^{\geq 0}$. Assume that Conjecture B holds true. Then $\mathbb{F}$ is the field of values of an irreducible character $\chi$ of some finite group with $\nu_{p}(\chi(1))=b$ if and only if
$\nu_{p}([\mathbb{F}_{p^{a}}:\mathbb{F}])\leq b.$
The next result presents evidence in support of Conjecture B.
THEOREM D.
Let $\chi$ be an irreducible character of some finite group $G$. Suppose $a:=\nu_{p}(c(\chi))\geq 1$. Then $\chi(1)$ is divisible by ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$ if one the following cases occurs.
  1. (a)
    $G$ is $p$-solvable;
  2. (b)
    $G$ is an alternating group;
  3. (c)
    $G$ is a general linear or general unitary group.
Next we turn to the related problem of classifying the fields of values of the irreducible characters of finite groups of order not divisible by $p$. This leads us to the following basic problem: is it always true that ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$? Interestingly, the recently proposed inductive Feit condition ([BKNT25]) implies what seemed to be an intractable problem.
THEOREM E.
Suppose that the finite simple groups satisfy the inductive Feit condition. Then the following hold:
  1. (a)
    If $G$ is a finite group and $\chi\in{\rm Irr}(G)$, then ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$.
  2. (b)
    Let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension and let $p$ be a prime. Then $\mathbb{F}={\mathbb{Q}}(\chi)$ for some $\chi\in{\rm Irr}(G)$, where $G$ is a finite group and $p$ does not divide $|G|$, if and only if both $c(\mathbb{F})$ and ${\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ are not divisible by $p$.
Theorems A, D, and E are proved in Sections 2, 3, and 4, respectively.
2. THEOREM A
Our notation is mostly standard and essentially follows [I76]. We begin by recalling a result of G. M. Cram that will be used frequently.
Lemma 2.1.
Let $G$ be a finite group and $N\trianglelefteq\,G$. Let $\chi\in\operatorname{Irr}(G)$ and $\theta$ be an irreducible constituent of $\chi_{N}$. Let $\psi$ be the Clifford correspondent of $\chi$ with respect to $\theta$. Then ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$ and ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$ divides $\chi(1)/\psi(1)$.
Proof.
This follows from [C88, Lemma 1.1]. ∎
The next lemma easily implies, using direct products, that if $\mathbb{F}$ is the field of values of an irreducible character $\chi$ with $\chi(1)_{p}=p^{b}$ for $b\in{\mathbb{Z}}_{\geq 0}$, then $\mathbb{F}$ also arises as the field of values of a character whose degree has any $p$-part at least $p^{b}$. Therefore, to classify the fields of values of irreducible characters whose degree has a prescribed $p$-part, the question we really have to solve is: given an abelian number field $\mathbb{F}$, what is the smallest possible $p$-part of the degree of an irreducible character $\chi$ such that ${\mathbb{Q}}(\chi)=\mathbb{F}$?
Lemma 2.2.
Let $d\in{\mathbb{Z}}_{\geq 1}$. Then there exists a solvable group $G$ and a rational-valued character $\chi\in\operatorname{Irr}(G)$ such that $\chi(1)=d$.
Proof.
This is [HT23, Lemma 2.1]. ∎
We also need the following.
Lemma 2.3.
Let $n\in{\mathbb{Z}}_{\geq 1}$ and $\mathbb{F}\supseteq{\mathbb{Q}}$ be a subfield of ${\mathbb{Q}}_{n}$. Then there exists an irreducible character $\chi$ (of some finite solvable group) such that ${\mathbb{Q}}(\chi)=\mathbb{F}$ and $\chi(1)=[{\mathbb{Q}}_{n}:\mathbb{F}]$.
Proof.
This is [FG72, Theorem 2.2]. See also [NT21, Theorem 2.2]. ∎
For solvable groups, Theorem C follows from known results.
Theorem 2.4.
Let $\mathbb{F}$ be an abelian extension of $\mathbb{Q}$} and $b\in{\mathbb{Z}}_{\geq 0}$. Then $\mathbb{F}$ is the field of values of an irreducible character $\chi$ of a finite solvable group with $\nu_{p}(\chi(1))=b$ if and only if
$\nu_{p}([{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b.$
Proof.
The ‘only if’ implication directly follows from Cram’s theorem ([C88, Theorem 0.1]). For the reverse implication, let $a:=\nu_{p}([{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b$. By Lemma 2.3, there exists a solvable group $H$ and a character $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\psi)=\mathbb{F}$ and $\psi(1)=[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$. On the other hand, by Lemma 2.2, there exists a solvable group $K$ and a rational-valued character $\varphi\in\operatorname{Irr}(K)$ such that $\varphi(1)=p^{b-a}$.
Let $G:=H\times K$ and $\chi:=\psi\otimes\varphi\in\operatorname{Irr}(G)$. We have ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\psi,\varphi)={\mathbb{Q}}(\psi)=\mathbb{F}$ and $\nu_{p}(\chi(1))=\nu_{p}(\psi(1)\varphi(1))=\nu_{p}(\psi(1))+\nu_{p}(\varphi(1))=a+(b-a)=b$. ∎
The next result we require is the $p^{\prime}$-degree case of Theorem A.
Theorem 2.5 ([NT21], Theorem B2).
Let $\mathbb{F}$ be an abelian extension of $\mathbb{Q}$} with $\nu_{p}(c(\mathbb{F}))=a$. Suppose that ${\mathbb{Q}}_{p^{a}}:(\mathbb{F}\cap{\mathbb{Q}}_{p^{a}})]$ is not divisible by $p$. Then there exists a finite group $G$ and $\chi\in\operatorname{Irr}_{p^{\prime}}(G)$ such that ${\mathbb{Q}}(\chi)=\mathbb{F}$.
One of the key constructions in producing the desired group and character in Theorem 2.5 is to reduce from a character field of values to a smaller one where the extension is cyclic, see [NT21, Theorem 3.3]. For our purpose, we need to generalize this to arbitrary (abelian) extensions.
Theorem 2.6.
Let $\chi$ be an irreducible character of a finite group $G$ and set ${\mathbb{E}}:={\mathbb{Q}}(\chi)$. Then, for any subfield $\mathbb{F}$ of $\mathbb{E}$}, there exists a finite group $H$ and a character $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\psi)=\mathbb{F}$ and $\psi(1)=n\chi(1)^{n}$, where $n:=[{\mathbb{E}}:\mathbb{F}]$.
Proof.
Write $\mathcal{G}:=\mathrm{Gal}({\mathbb{E}}/\mathbb{F})$ as a direct product ${\sf C}_{n_{1}}\times{\sf C}_{n_{2}}\times\cdots\times{\sf C}_{n_{k}}$ of cyclic groups of orders $n_{1},...,n_{k}$ with $n_{i}>1$ and $n_{i}$ is divisible by $n_{i+1}$. Note that $n:=\prod_{i=1}^{k}n_{i}=|\mathcal{G}|$. Let $\tau_{i}$ be a generator of ${\sf C}_{i}$, so that
$\mathcal{G}=\{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}\mid 1\leq a_{i}\leq n_{i}\text{ for every }1\leq i\leq k\}$
Consider the $n$-iterated direct product of $G$ and view its elements as $k$-dimensional matrices:
$G^{n}:=\{X=(x_{a_{1},...,a_{k}})_{1\leq a_{i}\leq n_{i}\text{ for every }1\leq i\leq k}\mid x_{a_{1},...,a_{k}}\in G\}.$
Let $\tau_{i}$ act on $G^{n}$ by permuting the $i$th-dimensional indices of the entries $x_{a_{1},...,a_{k}}$ but fixing the indices at the other dimensions, in the following way
$(x_{a_{1},...,a_{k}})^{\tau_{i}}=(x_{a_{1},...,a_{i-1},a_{i}-1,a_{i+1},...,a_{k}}),$
where two indices at the $i$th-dimension are identified if they have the same residue modulo $n_{i}$. This action induces an embedding of $\mathcal{G}={\sf C}_{n_{1}}\times{\sf C}_{n_{2}}\times\cdots\times{\sf C}_{n_{k}}$ into the symmetric group ${\mathrm{Sym}}_{n}$. Let
$H:=G^{n}\rtimes\mathcal{G}$
be the corresponding semidirect product.
Let
$\varphi:=\bigotimes_{1\leq a_{i}\leq n_{i}}\varphi_{a_{1},...,a_{k}}\in\operatorname{Irr}(G^{n}),$
where
$\varphi_{a_{1},...,a_{k}}=\chi^{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}}\in\operatorname{Irr}(G);$
that is,
$\varphi(X)=\prod_{1\leq a_{i}\leq n_{i}\atop 1\leq i\leq k}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...,a_{k}}).$
The action of a typical element $\tau=\tau_{1}^{b_{1}}\cdots\tau_{k}^{b_{k}}$ of $\mathcal{G}$ on $\varphi$ is given by
$(\varphi^{\tau})_{a_{1},...,a_{k}}=\chi^{\tau_{1}^{a_{1}+b_{1}}\cdots\tau_{k}^{a_{k}+b_{k}}}.$
It is clear that $\varphi^{\tau}\neq\varphi$ whenever $\tau\neq 1$. Therefore, the stabilizer $I_{H}(\varphi)$ of $\varphi$ in $H$ is simply $G^{n}$, which in turn implies that
$\psi:=\varphi^{H}\in\operatorname{Irr}(H).$
We have $\psi(1)=[H:G^{n}]\varphi(1)=n\chi(1)^{n}$. The rest of the proof will establish that ${\mathbb{Q}}(\psi)=\mathbb{F}$. This can be done by an suitable modification of the arguments in the proof of [NT21, Theorem 3.3].
By the definition of $\varphi$, we know ${\mathbb{Q}}(\varphi)={\mathbb{Q}}(\chi)={\mathbb{E}}$. The character-induction formula then implies that ${\mathbb{Q}}(\psi)\subseteq{\mathbb{E}}$. Furthermore, for each $X\in G^{n}$ and $j\in\{1,...,k\}$,
$\varphi(X^{\tau_{j}})=\prod_{1\leq a_{i}\leq n_{i}\atop 1\leq i\leq k}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...a_{j-1},a_{j}-1,a_{j+1},...,a_{k}}).$
On the other hand,
$\varphi^{\tau_{j}}(X)=\prod_{1\leq a_{i}\leq n_{i}\atop 1\leq i\leq k}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{j-1}^{a_{j-1}}\tau_{j}^{a_{j}+1}\tau_{j+1}^{a_{j+1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...,a_{k}}).$
If all the $a_{1},...,a_{j-1},a_{j+1},...,a_{k}$ are fixed, the sub-products along the $j$th dimension are equal:
$\prod_{1\leq a_{j}\leq n_{j}}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...a_{j-1},a_{j}-1,a_{j+1},...,a_{k}})=\prod_{1\leq a_{j}\leq n_{j}}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{j-1}^{a_{j-1}}\tau_{j}^{a_{j}+1}\tau_{j+1}^{a_{j+1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...,a_{k}})$
It follows that
$\varphi(X^{\tau_{j}})=\varphi^{\tau_{j}}(X).$
Since $\mathcal{G}$ is generated by $\tau_{j},1\leq j\leq k$, we conclude that
$\varphi(X^{\tau})=\varphi^{\tau}(X)$
for every $\tau\in\mathcal{G}$. That is, the Galois action of $\mathcal{G}$ on $\varphi$ is compatible with our defined action of $\mathcal{G}$ on $G^{n}$.
Now, for every $X\in G^{n}$ and $\tau\in\mathcal{G}$, by the Clifford theorem, we have
$\psi(X)^{\tau}=$ $\left(\sum_{\tau^{\prime}\in\mathcal{G}}\varphi(X^{\tau^{\prime}})\right)^{\tau}=\left(\sum_{\tau^{\prime}\in\mathcal{G}}\varphi(X)^{\tau^{\prime}}\right)^{\tau}$
$=$ $\sum_{\tau^{\prime}\in\mathcal{G}}\varphi(X)^{\tau\tau^{\prime}}=\sum_{\tau^{\prime}\in\mathcal{G}}\varphi(X)^{\tau^{\prime}}=\psi(X).$
Therefore, $\psi(X)$ is fixed by $\mathcal{G}$ point-wisely. As $\psi$ vanishes outside $G^{n}$, we deduce that ${\mathbb{Q}}(\psi)={\mathbb{Q}}(\psi_{G^{n}})$ is contained in the fixed field of $\mathcal{G}$ in $\mathbb{E}$}. That is, ${\mathbb{Q}}(\psi)\subseteq\mathbb{F}$.
By Lemma 2.1, we have ${\mathbb{Q}}(\varphi):{\mathbb{Q}}(\psi)]$ divides $H:G^{n}]=n$. As mentioned already that ${\mathbb{Q}}(\varphi)={\mathbb{E}}$, it follows that ${\mathbb{E}}:{\mathbb{Q}}(\psi)]\leq n=[{\mathbb{E}}:\mathbb{F}]$. This and the conclusion of the preceding paragraph imply that ${\mathbb{Q}}(\psi)=\mathbb{F}$. The proof is complete. ∎
We can now complete the proof of Theorem A, which is restated.
Theorem 2.7.
Let $\mathbb{F}$ be an abelian extension of $\mathbb{Q}$} and $a:=\nu_{p}(c(\mathbb{F}))$. Let $c:=\nu_{p}([\mathbb{F}_{p^{a}}:\mathbb{F}])$. Then, for every integer $b\geq c$, there exists a finite group $G$ and a character $\chi\in\operatorname{Irr}(G)$ such that ${\mathbb{Q}}(\chi)=\mathbb{F}$ and $\nu_{p}(\chi(1))=b$.
Proof.
Let ${\mathbb{E}}:={\mathbb{Q}}_{p^{a}}(\mathbb{F})$. It is clear that $\nu_{p}(c({\mathbb{E}}))=a$ and ${\mathbb{Q}}_{p^{a}}:({\mathbb{E}}\cap{\mathbb{Q}}_{p^{a}})]=1$. Using Theorem 2.5, we obtain a finite group $H$ and an irreducible character $\psi$ of $H$ of $p^{\prime}$-degree such that ${\mathbb{Q}}(\psi)={\mathbb{E}}$.
By the hypothesis, we have $\mathbb{F}_{p^{a}}:\mathbb{F}]=[{\mathbb{E}}:\mathbb{F}]=p^{c}f$ for some $f\in{\mathbb{Z}}^{+}$ not divisible by $p$. Now applying Theorem 2.6 to the extension ${\mathbb{E}}/\mathbb{F}$, we then obtain a finite group $M$ and $\varphi\in\operatorname{Irr}(M)$ such that
$\varphi(1)=p^{c}f\psi(1)^{p^{c}f}\text{ and }{\mathbb{Q}}(\varphi)=\mathbb{F}.$
Next, we use Lemma 2.2 to have a finite group $N$ and some rational-valued character $\lambda\in\operatorname{Irr}(N)$ such that $\lambda(1)=p^{b-c}$.
Finally, let $G:=M\times N$ and $\chi:=\varphi\otimes\lambda$ – an irreducible character of $G$. This $\chi$ is indeed the required character. To see that, we have
$\chi(1)=\varphi(1)\lambda(1)=p^{c}f\psi(1)^{p^{c}f}\cdot p^{b-c}=p^{b}f\psi(1)^{p^{c}f}.$
Recall that both $f$ and $\psi(1)$ are not divisible by $p$. Therefore $\nu_{p}(\chi(1))=b$. Moreover,
${\mathbb{Q}}(\chi)={\mathbb{Q}}(\varphi,\lambda)={\mathbb{Q}}(\varphi)=\mathbb{F},$
as desired. ∎
3. Theorem D
We begin with the following lemma about $p$-groups.
Lemma 3.1.
Suppose that $P$ is a $p$-group. Let $\alpha\in{\rm Irr}(P)$ be with $c(\alpha)=p^{a}$ and $a\geq 1$.
  1. (a)
    If $p$ is odd or $p^{a}=4$, then ${\mathbb{Q}}_{p^{a}}(\alpha)={\mathbb{Q}}_{p}(\alpha)$.
  2. (b)
    If $p=2$ and $p^{a}\geq 8$, then $|{\mathbb{Q}}_{p^{a}}(\alpha):{\mathbb{Q}}_{p}(\alpha)|\leq 2$ divides $\alpha(1)$.
Proof.
If $p$ is odd, then we know that ${\mathbb{Q}}(\alpha)={\mathbb{Q}}_{p^{a}}$ by [NT21, Theorem 2.3]. If $p=2$ and $a=2$, then ${\mathbb{Q}}(\alpha)\subseteq{\mathbb{Q}}(i)$ and necessarily ${\mathbb{Q}}(\alpha)={\mathbb{Q}}(i)$, so there is nothing to prove. Assume now that $p=2$ and $a\geq 3$. By elementary Galois theory, we know that ${\mathbb{Q}}_{2^{a}}/{\mathbb{Q}}_{8}$ is a cyclic extension with intermediate fields ${\mathbb{Q}}_{8}\subseteq{\mathbb{Q}}_{16}\subseteq\ldots\subseteq{\mathbb{Q}}_{2^{a}}$. Therefore ${\mathbb{Q}}_{8}(\alpha)={\mathbb{Q}}_{2^{a}}={\mathbb{Q}}_{2^{a}}(\alpha)$. Also ${\mathbb{Q}}_{8}\cap{\mathbb{Q}}(\alpha)$ cannot be $\mathbb{Q}$}, since ${\mathbb{Q}}_{2^{a}}$ has only three subfields of degree 2 and all of them are contained in ${\mathbb{Q}}_{8}$. If $\alpha$ is linear, notice that ${\mathbb{Q}}_{p^{a}}(\alpha):{\mathbb{Q}}_{p}(\alpha)]=1$. Part (b) easily follows. ∎
Theorem 3.2.
Let $G$ be a $p$-solvable finite group, and let $\chi\in{\rm Irr}(G)$ be with $a=\nu_{p}(c(\chi))\geq 1$. Then ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$ divides $\chi(1)$.
Proof.
Notice that ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$ is a power of $p$. We argue by induction on $|G|$. Suppose that there is $N\trianglelefteq\,G$, and $\eta^{G}=\chi$, where $\eta\in{\rm Irr}(V)$ is a Clifford correspondent of $\chi$. By Lemma 2.1, we know that ${\mathbb{Q}}(\eta):{\mathbb{Q}}(\chi)]$ divides $|G:V|$. We have that $c(\chi)$ divides $c(\eta)$. Write $c(\eta)=p^{b}$, where $b\geq 1$. By induction, we have that ${\mathbb{Q}}_{p^{a}}(\eta):{\mathbb{Q}}_{p}(\eta)]$ divides ${\mathbb{Q}}_{p^{b}}(\eta):{\mathbb{Q}}_{p}(\eta)]$ divides $\eta(1)$. If $K={\mathbb{Q}}_{p^{a}}(\chi)\cap{\mathbb{Q}}_{p}(\eta)$, by elementary Galois theory, we have that $K:{\mathbb{Q}}_{p}(\chi)]$ divides ${\mathbb{Q}}(\eta):{\mathbb{Q}}(\chi)]$ which divides $|G:V|$, and we are done.
We may now assume that $\chi$ is quasi-primitive. Hence, $\chi=\alpha\beta$, where $\alpha$ is $p$-special and $\beta$ is $p^{\prime}$-special (see [I18, Theorem 2.9]). We claim that ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]=1$, in this case. First, we prove that ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\alpha,\beta)$. Certainly, ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\alpha,\beta)$. However, if $\sigma\in{\rm Gal}({\mathbb{Q}}(\alpha,\beta)/{\mathbb{Q}}(\chi)$, then we obtain that $\alpha^{\sigma}=\alpha$ and $\beta^{\sigma}=\beta$, by the uniqueness of the product decomposition. Since ${\mathbb{Q}}(\alpha)={\mathbb{Q}}(\alpha_{P})$, where $P\in{\rm Syl}_{p}(G)$ (by the uniqueness of the restriction from $p$-special characters), we have that $c(\alpha\beta)=c(\alpha)c(\beta)$. If $p$ is odd or $p^{a}=4$, then ${\mathbb{Q}}_{p^{a}}={\mathbb{Q}}_{p}(\alpha)$, by Lemma 3.1(a). Then ${\mathbb{Q}}_{p^{a}}(\chi)={\mathbb{Q}}_{p^{a}}(\alpha,\beta)={\mathbb{Q}}_{p^{a}}(\beta)$ and ${\mathbb{Q}}_{p}(\chi)={\mathbb{Q}}_{p}(\alpha,\beta)={\mathbb{Q}}_{p^{a}}(\beta)$. Suppose finally that $p=2$ and $a\geq 3$. If $\chi$ has odd-degree, then the theorem follows from [NT21, Theorem A1]. So we may assume that $\chi(1)$ is even. By Lemma 3.1, we have that ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]\leq 2$, and the theorem follows. ∎
Theorem 3.3.
Let $n\geq 1$ and $\chi\in{\rm Irr}({\mathrm{Alt}}_{n})\cup{\rm Irr}({\mathrm{Sym}}_{n})$. Then $\nu_{p}(c(\chi))\leq 1$. In particular, Theorem D holds for the symmetric and alternating groups.
Proof.
It is well known that the characters of ${\mathrm{Sym}}_{n}$ are all rational, so the result is immediate in this case. Hence, we may assume that $\chi\in\operatorname{Irr}({\mathrm{Alt}}_{n})$. For background on the ordinary irreducible characters of ${\mathrm{Alt}}_{n}$, we refer the reader to [JK81, §5].
Suppose that $\Psi$ is the irreducible character of ${\mathrm{Sym}}_{n}$ lying over $\chi$, and let $\lambda$ be the partition of $n$ labeling $\Psi$. If $\lambda$ is not self-conjugate, then $\chi=\Psi_{{\mathrm{Alt}}_{n}}$ and the result is again trivial. Thus we assume that $\lambda$ is self-conjugate. In this case, $\Psi_{{\mathrm{Alt}}_{n}}=\chi+\chi^{\sigma},$ where $\sigma$ is an odd permutation in ${\mathrm{Sym}}_{n}$.
Let $h(\lambda)=(h_{11}^{\lambda},\dots,h_{kk}^{\lambda})$ denote the partition of $n$ whose parts $h_{ii}^{\lambda}$ are the hook lengths at the diagonal nodes $(i,i)$ of the Young diagram of $\lambda$. Here $k$ is the length of the main diagonal. The conjugacy class of ${\mathrm{Sym}}_{n}$ of cycle type $h(\lambda)$ splits into two conjugacy classes of ${\mathrm{Alt}}_{n}$, say $h(\lambda)^{+}$ and $h(\lambda)^{-}$.
According to [JK81, Theorem 2.5.13], the values of $\chi$ (and of $\chi^{\sigma}$) are rational at all elements, except possibly at those belonging to $h(\lambda)^{\pm}$. In fact, one has
${\mathbb{Q}}(\chi)={\mathbb{Q}}\!\left(\sqrt{(-1)^{(n-k)/2}\,h_{11}^{\lambda}\cdots h_{kk}^{\lambda}}\right).$
Let $d$ denote the square-free part of $h_{11}^{\lambda}\cdots h_{kk}^{\lambda}$, that is, the product of the prime divisors occurring with odd exponents. Note that all the hook lengths $h_{ii}^{\lambda}$ are odd and that
$n-k=\sum_{i=1}^{k}(h_{ii}^{\lambda}-1)$
is even. We shall use $c(z)$ to denote the conductor of a sum of roots of unity $z$, which is the smallest positive integer $t$ such that $z\in{\mathbb{Q}}_{t}$. If $n-k\equiv 0\pmod{4}$, then $h_{11}^{\lambda}\cdots h_{kk}^{\lambda}\equiv 1\pmod{4}$, and hence
$c(\chi)=c\!\left(\sqrt{h_{11}^{\lambda}\cdots h_{kk}^{\lambda}}\right)=c(\sqrt{d})=d.$
On the other hand, if $n-k\equiv 2\pmod{4}$, then $h_{11}^{\lambda}\cdots h_{kk}^{\lambda}\equiv 3\pmod{4}$, and hence
$c(\chi)=c\!\left(\sqrt{-h_{11}^{\lambda}\cdots h_{kk}^{\lambda}}\right)=c(\sqrt{-d})=d.$
It follows that $\nu_{p}(c(\chi))=1$ if $p$ divides $d$, and $\nu_{p}(c(\chi))=0$ otherwise. This completes the proof. ∎
We now prove Theorem D for finite general linear groups $G=\operatorname{GL}_{n}(q)$, where $n\geq 1$ and $q$ is a prime power, with a natural module $V=\mathbb{F}_{q}^{n}=\langle e_{1},\ldots,e_{n}\rangle_{\mathbb{F}_{q}}$. It is convenient for us to use the Dipper-James classification of complex irreducible characters of $G$, as described in [J86]. Namely, every $\chi\in\operatorname{Irr}(G)$ can be written uniquely, up to a permutation of the pairs $\{(s_{1},\lambda_{1}),\dots,(s_{m},\lambda_{m})\}$, in the form
(3.1) $\chi=S(s_{1},\lambda_{1})\circ S(s_{2},\lambda_{2})\circ\ldots\circ S(s_{m},\lambda_{m}).$
Here, $s_{i}\in\overline{\mathbb{F}_{q}}^{\times}$ has degree $d_{i}$ over $\mathbb{F}_{q}$, $\lambda_{i}\vdash k_{i}$, $\sum^{m}_{i=1}k_{i}d_{i}=n$, and the $m$ elements $s_{i}$ have pairwise distinct minimal polynomials over $\mathbb{F}_{q}$. In particular, $S(s_{i},\lambda_{i})$ is an irreducible character of $\operatorname{GL}_{k_{i}d_{i}}(q)$. Furthermore, there is a parabolic subgroup $P_{\chi}=U_{\chi}L_{\chi}$ of $G$ with Levi subgroup $L_{\chi}=\operatorname{GL}_{k_{1}d_{1}}(q)\times\ldots\times\operatorname{GL}_{k_{m}d_{m}}(q)$ and unipotent radical $U_{\chi}$. The (outer) tensor product
$\psi:=S(s_{1},\lambda_{1})\otimes S(s_{2},\lambda_{2})\otimes\ldots\otimes S(s_{m},\lambda_{m})$
is an $L_{\chi}$-character, and $\chi$ is obtained from $\psi$ via the Harish-Chandra induction $R^{G}_{L_{\chi}}$, i.e. we first inflate $\psi$ to a $P_{\chi}$-character and then induce it to $G$.
Lemma 3.4.
In the above notation, assume that the integer $N\geq 1$ is divisible by the order of every $s_{i}$, $1\leq i\leq m$, and that $\sigma\in\mathrm{Gal}({\mathbb{Q}}_{N}/{\mathbb{Q}})$ sends a primitive $N^{\mathrm{th}}$ root of unity $\zeta\in{\mathbb{C}}^{\times}$ to $\zeta^{k}$, $1\leq k<N$. Then for the character $\chi$ in (3.1),
$\sigma(\chi)=S(s_{1}^{k},\lambda_{1})\circ S(s_{2}^{k},\lambda_{2})\circ\ldots\circ S(s_{m}^{k},\lambda_{m}).$
Proof.
First we show that $s_{i}^{k}$ also has degree $d_{i}$ over $\mathbb{F}_{q}$, for any $i$. Indeed, $s_{i}\in\mathbb{F}_{q^{d_{i}}}$ implies that $s_{i}^{k}\in\mathbb{F}_{q^{d_{i}}}$, and hence $s_{i}^{k}$ has degree $d^{\prime}|d_{i}$. Conversely, as $\gcd(k,N)=1$, we can find $l\in{\mathbb{Z}}_{\geq 1}$ such that N|(kl1)N|(kl-1). Now $s_{i}=s_{i}^{kl}\in\mathbb{F}_{q^{d^{\prime}}}$, and so $d_{1}|d^{\prime}$, whence $d^{\prime}=d_{i}$.
Since the Harish-Chandra induction $R^{G}_{L_{\chi}}$ commutes with Galois action, it suffices to prove the claim in the case $m=1$. Now, if $d_{1}>1$, then $\chi=S(s_{1},\lambda_{1})$ is Lusztig induced from the character $S(s_{1},\lambda_{1})$ of the Levi subgroup $L=\operatorname{GL}_{k_{1}}(q^{d_{1}})$ of $G=\operatorname{GL}_{k_{1}d_{1}}(q)$, see [FS82, p. 116]. Since the Lefschetz numbers are all integers (see e.g. [DM91, p. 140]), Green functions involved in the Lusztig induction $R^{G}_{L}$ rational-valued, and thus $R^{G}_{L}$ also commutes with Galois actions. Thus it suffices to prove the claim in the case $m=d_{1}=1$. In the latter case, $\chi=S(1,\lambda_{1})\hat{s_{1}}$, where $S(1,\lambda_{1})$ is a rational (unipotent) character, and $\hat{s_{1}}$ is a linear character of $G/[G,G]$ of order equal to the order of $s_{1}$, see [FS82, (1.16)], so we are done. ∎
Lemma 3.5.
Let $p$ be a prime and $A={\sf C}_{p^{n}}\times{\sf C}_{p}$ for some $n\geq 1$. If $B$ is a subgroup of $A$, then there is some $1\leq m\leq n$ such that $B\cong\mathsf{C}_{p^{m}}$ or $B\cong\mathsf{C}_{p^{m}}\times\mathsf{C}_{p}$.
Proof.
Write $A=A_{1}\times A_{2}$ with $A_{1}=\langle x\rangle\cong\mathsf{C}_{p^{n}}$ and $A_{2}=\langle y\rangle\cong\mathsf{C}_{p}$, and consider the natural projection $\pi:A\to A_{1}$ with kernel $A_{2}$. If $\pi|_{B}$ is injective then $B$ is cyclic of order $p^{m}$ with $m\leq n$, and we are done. Otherwise $1\neq\operatorname{Ker}(\pi|_{B})$, so $y\in B$, whence $B=A_{2}\times(A_{1}\cap B)$, and we are done again. ∎
Theorem 3.6.
Theorem D holds for finite general linear groups $G=\operatorname{GL}_{n}(q)$ and finite general unitary groups $\operatorname{GU}_{n}(q)$.
Proof.
Since the statement is trivial when $a=1$ or $n=1$, we may assume that $a,n\geq 2$.
The case of $\operatorname{GU}_{n}(q)$ is entirely similar to that of $\operatorname{GL}_{n}(q)$, changing only $q$ to q-q (in character degrees and subgroup orders, as well as in the definition of $[t]$ for $t\in\overline{\mathbb{F}_{q}}^{\times}$) and Harish-Chandra induction to Lusztig induction. So we will provide the details only for $G=\operatorname{GL}_{n}(q)$. Note that $p\nmid q$ since (3.1) shows that $c(\chi)$ divides $N:=\mathrm{lcm}(o(s_{1}),\ldots,o(s_{m}))$ which divides $\prod^{n}_{i=1}(q^{i}-1)$. So $q$ is power of some prime $r\neq p$.
(a) Here we show that if a $p$-element $\sigma\in\mathrm{Gal}({\mathbb{Q}}_{N}/{\mathbb{Q}}_{p}(\chi))$ induces an element of $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$ of order $p^{a_{1}}$ for some $a_{1}\geq 0$, then $\chi(1)$ is divisible by $p^{a_{1}}$, in fact by $2^{a_{1}+1}$ if $p=2\leq a_{1}$.
The case $a_{1}=0$ is trivial. so we may assume $a_{1}\geq 1$, and note that $\chi$ is $\sigma$-invariant. Writing $\chi$ in the form (3.1) and applying Lemma 3.4 with $\sigma(\zeta)=\zeta^{f}$, we see that the induced map $\tau:S(s,\lambda)\mapsto S(s^{f},\lambda)$ acts on the set $S$ of the characters $S(s_{i},\lambda_{i})$, $1\leq i\leq m$.
(a1) First we show that if some $\tau$-orbit $\mathcal{O}$} on $S$ has length $p^{b}$ with $b\geq 0$, say $S(s_{1},\lambda_{1})\in{\mathcal{O}}$, and $p$ divides $q^{k}-1$ with $k:=k_{1}d_{1}$, then both
$N_{1}:=\frac{|\operatorname{GL}_{kp^{b}}(q)|_{r^{\prime}}}{|\operatorname{GL}_{k}(q)|_{r^{\prime}}^{p^{b}}}$
and $\chi(1)$ are divisible by $p^{b}$, in fact by $p^{b+1}$ if $p=2\leq b$.
Indeed, the statement is obvious if $b=0$. Assume $b\geq 1$. Then for any $1\leq i\leq b$, $p|(q^{kp^{i}}-1)/(q^{kp^{i-1}}-1)$, hence $p^{b}$ divides $(q^{kp^{b}}-1)/(q^{k}-1)$. Since
$N_{1}:=N_{2}\cdot\frac{q^{kp^{b}}-1}{q^{k}-1},\mbox{ where }N_{2}:=\frac{|\operatorname{GL}_{kp^{b}-1}(q)|_{r^{\prime}}}{|\operatorname{GL}_{k}(q)|_{r^{\prime}}^{p^{b}-1}|\operatorname{GL}_{k-1}(q)|_{r^{\prime}}}$
is an integer, $p^{b}|N_{1}$. Next, $N_{1}$ divides
$\frac{|\operatorname{GL}_{n}(q)|_{r^{\prime}}}{\prod^{m}_{i=1}|\operatorname{GL}_{k_{i}d_{i}}(q)|_{r^{\prime}}}.$
So $N_{1}|\chi(1)$, and hence $p^{b}|\chi(1)$.
Assume in addition that $p=2\leq b$. Then we can express $N_{2}$ as
$\biggl(\frac{|\operatorname{GL}_{k\cdot 2^{b-1}-1}(q)|_{r^{\prime}}}{|\operatorname{GL}_{k}(q)|_{r^{\prime}}^{2^{b-1}-1}|\operatorname{GL}_{k-1}(q)|_{r^{\prime}}}\biggr)^{2}\cdot\frac{|\operatorname{GL}_{k\cdot 2^{b}-1}(q)|_{r^{\prime}}}{|\operatorname{GL}_{k\cdot 2^{b-1}}(q)|_{r^{\prime}}|\operatorname{GL}_{k\cdot 2^{b-1}-1}(q)|_{r^{\prime}}}\cdot\frac{q^{k\cdot 2^{b-1}}-1}{q^{k}-1},$
where the first two factors are integers, and the third one is an even integer. So $2^{b+1}$ divides both $N_{1}$ and $\chi(1)$.
(a2) For any $t\in\overline{\mathbb{F}_{q}}^{\times}$ of degree say $l$ over $\mathbb{F}_{q}$, let $[t]$ denote the set $\{t,t^{q},\ldots,t^{q^{l-1}}\}$. Suppose that some $\tau$-orbit $\mathcal{O}$} on $S$ has length $p^{b}$ with $b\geq 0$, say $S(s_{1},\lambda_{1})\in{\mathcal{O}}$, and $p|(q^{d}-1)$ with $d:=d_{1}$. As $\tau^{p^{b}}$ fixes $S(s_{1},\lambda_{1})$, $\tau^{p^{b}}$ acts on the set $s_{1}]$. Suppose in addition that some $\tau^{p^{b}}$-orbit on $s_{1}]$ has length $p^{c}$ for some $c\geq 0$. Then we show that $p^{b+c}|\chi(1)$, and in fact $2^{b+c+1}|\chi(1)$ if $p=2\leq b+c$.
Indeed, write $k:=k_{1}$. If $c=0$, then we are done by (a1). So we assume that $c\geq 1$. By (a1),
$N_{1}:=\frac{|\operatorname{GL}_{kdp^{b}}(q)|_{r^{\prime}}}{|\operatorname{GL}_{kd}(q)|_{r^{\prime}}^{p^{b}}}$
is divisible by $p^{b}$, in fact by $p^{b+1}$ if $p=2\leq b$.
Since the maps $\tau^{p^{b}}:x\mapsto x^{f^{p^{b}}}$ and $x\mapsto x^{q}$ have commuting actions on $s_{1}]$, with the latter acting transitively, we see that $p^{c}$ divides the size of any $\tau^{p^{b}}$-orbit on $s_{1}]$, which has cardinality $d$. Thus $p^{c}|d$; write $d=p^{c}d_{1}$. As $p\nmid q$, we have $q^{d}\equiv q^{d_{1}}\pmod{p}$, and so $p|(q^{d_{1}}-1)$. Hence $p$ divides each $q^{id_{1}}-1$ with $1\leq i\leq p^{c}-1$, and so $p^{p^{c}-1}$ divides $(q-1)(q^{2}-1)\ldots(q^{d-1}-1)$, which is a divisor of
$\frac{(q-1)(q^{2}-1)\ldots(q^{kd}-1)}{(q^{d}-1)(q^{2d}-1)\ldots(q^{kd}-1)}.$
The latter divides the degree of $S(s_{1},\lambda_{1})$, see [KT, Lemma 5.7(ii)]. Also, $c\geq 1$ implies that $p^{c}-1\geq 2^{c}-1\geq c$. It follows that the degree of $S(s_{1},\lambda_{1})$ is divisible by $p^{c}$, in fact by $p^{c+1}$ if $p=2\leq c$.
Thus $p^{b+c}|\chi(1)$; moreover $p^{b+c+1}|\chi(1)$ if $p=2\leq\max(b,c)$. It remains to consider the case $p=2$ and $b=c=1$. In this case $\chi(1)$ is divisible by $N_{1}\bigl(\deg(S(s_{1},\lambda_{1}))\bigr)^{2}$, so by $2^{3}$ as stated.
(a3) Now, if $\tau$ has some orbit of length $p^{b}$ on $S$ containing $S(s_{j},\lambda_{j})$ with $p|o(s_{j})$ and $\tau^{p^{b}}$ has some orbit of length $p^{c}$ on the set $s_{j}]$ with $b+c\geq a_{1}$, then we are done by (a2). So we may assume that $\tau_{1}:=\tau^{p^{a_{1}-1}}$ fixes every $s_{i}$ with $p|o(s_{i})$. Thus if $\tau_{1}(x)=x^{f_{1}}$, then $f_{1}-1$ is divisible by $o(s_{i})$ whenever $p|o(s_{i})$. On the other hand, (3.1) shows that $c(\chi)$ divides $\mathrm{lcm}(o(s_{1}),\ldots,o(s_{m}))$. It follows that $p^{a}=c(\chi)_{p}$ divides $f_{1}-1$. But in this case, $\tau_{1}$ acts trivially on both ${\mathbb{Q}}_{p^{a}}$ and ${\mathbb{Q}}(\chi)$, contrary to the assumption that $\tau_{1}$ has order $p$ in $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$.
(b) By [W09, Theorem 3.4.3], the restriction to ${\mathbb{Q}}_{p^{a}}$ gives an injective homomorphism $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))\to\mathrm{Gal}({\mathbb{Q}}_{p^{a}}/{\mathbb{Q}}_{p})$, and the latter is isomorphic to $\mathsf{C}_{p^{a-1}}$ if $p>2$ and to $\mathsf{C}_{2^{a-2}}\times\mathsf{C}_{2}$ if $p=2$. It follows that $|\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))|=p^{e}$ with $e\leq a-1$.
It suffices to prove the statement in the case $e\geq 1$. Now, if $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$ is cyclic, then we can find an element of order $p^{e}$ in $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$, and so $p^{e}|\chi(1)$ by (a), and we are done.
By Lemma 3.5, it remains to consider the case $p=2\leq e$ and $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))\cong\mathsf{C}_{2^{e-1}}\times\mathsf{C}_{2}$. Now, if $e\geq 3$, then then we can find an element of order $2^{e-1}\geq 4$ in $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$, and so $2^{e+1}|\chi(1)$ by (a), and we are done again. Suppose now that $e=2$ but $4\nmid\chi(1)$. The analysis in (a) then shows that, for every $\tau$ induced by a nontrivial element in $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$, $\tau$ has exactly one orbit of length $2$ on
$S^{\prime}:=\{(s_{i}^{q^{j}},\lambda_{i})\mid 1\leq i\leq m,0\leq j\leq d_{j}-1\},$
with all other orbits having length $1$. However, since $4\nmid\chi(1)$, the same holds for the induced action of $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))=\langle\sigma_{1},\sigma_{2}\rangle\cong\mathsf{C}_{2}^{2}$. Thus if $\tau_{i}$ denotes the induced action of $\sigma_{i}$, $i=1,2$, then we may assume that both $\tau_{1,2}$ act as the same $2$-cycle on $S^{\prime}$. But then the automorphism $\sigma_{1}\sigma_{2}$ of order $2$ acts trivially on $S^{\prime}$, a contradiction. ∎
Theorem D follows from Theorems 3.2, 3.3, and 3.6.
4. Theorem E
For the definition of the inductive Feit condition, we refer the reader to [BKNT25, §3].
Theorem 4.1.
Let $G$ be a finite group and $\chi\in\operatorname{Irr}(G)$ be non-linear. Assume that simple groups involved in $G$ satisfy the inductive Feit condition. Then there exist $H<G$ and $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$ and ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$ divides $|G:H|$.
Proof.
Suppose that $N\trianglelefteq\,G$, $\theta\in{\rm Irr}(N)$ is an irreducible constituent of $\chi_{N}$ and let $\eta\in{\rm Irr}(T|\theta)$ be the Clifford correspondent of $\chi$ over $\theta$. By Lemma 2.1, we have that ${\mathbb{Q}}(\eta):{\mathbb{Q}}(\chi)]$ divides $|G:T|$. If $T<G$, then we are done. Hence, we may assume that $\chi$ is quasi-primitive. By [BKNT25, Theorem 4.7], there exists $H<G$ such that ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\psi)$, and we are done. (Notice that this theorem is stated for primitive characters, but as written after, it also holds for quasi-primitive characters.) ∎
The following is Theorem E.
Corollary 4.2.
Suppose that the finite simple groups satisfy the inductive Feit condition. Then the following hold:
  1. (a)
    If $G$ is a finite group and $\chi\in{\rm Irr}(G)$, then ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$.
  2. (b)
    Let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension and let $p$ be a prime. Then $\mathbb{F}={\mathbb{Q}}(\chi)$ for some $\chi\in{\rm Irr}(G)$, where $G$ is a finite group and $p$ does not divide $|G|$, if and only if both $c(\mathbb{F})$ and ${\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ are not divisible by $p$.
Proof.
We argue by induction on $|G|$. If $\chi$ is linear, then ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]=1$, and we are done. We may assume that $\chi$ is not linear. By Theorem 4.1, there exists $H<G$ and $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$ and ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$ divides $|G:H|$. By induction, we have that ${\mathbb{Q}}_{c(\psi)}:{\mathbb{Q}}(\psi)]$ divides $|H|$. Since ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$, we have that $c(\chi)$ divides $c(\psi)$. Let $K={\mathbb{Q}}_{c(\chi)}\cap{\mathbb{Q}}(\psi)$. By elementary Galois theory, ${\mathbb{Q}}_{c(\chi)}:K]$ divides ${\mathbb{Q}}_{c(\psi)}:{\mathbb{Q}}(\psi)]$, which divides $|H|$. Also, $K:{\mathbb{Q}}(\chi)]$ divides ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$, which divides $|G:H|$, and part (a) follows.
Now, let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension and let $p$ be a prime. By Lemma 2.3, there exists a group $G$ of order $c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ and $\chi\in\operatorname{Irr}(G)$ such that ${\mathbb{Q}}(\chi)=F$. Hence, if $c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ is not divisible by $p$, we have that $G$ is a $p^{\prime}$-group. Conversely, assume that $G$ is a $p^{\prime}$-group and $\chi\in{\rm Irr}(G)$. Since $c(\chi)$ divides $|G|$, we have that $c(\chi)$ is not divisible by $p$. By part (a), we have that ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$, and we are done. ∎
For solvable groups, it was proven in [C88, Theorem 0.1] that in fact ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $\chi(1)$, but this is not true in general (as shown, for instance, by ${\mathrm{Alt}}_{5}$ and the irreducible characters of degree 3).