### v2

# Primes and The Field of Values of Characters

Abstract.
Let $p$ be a prime. For $p=2$, the fields of values of the complex irreducible characters of finite groups whose degrees are not divisible by $p$ have been classified; for odd primes $p$, a conjectural classification has been proposed. In this work, we extend this conjecture to characters whose degrees are divisible by arbitrary powers of $p$, and we provide some evidence supporting its validity.
Key words and phrases:
Character, field of values, conductor
2020 Mathematics Subject Classification:
Primary 20C15; Secondary 20C30, 20C33
The first author gratefully acknowledges the support of the AMS-Simons Research Enhancement Grant (AWD-000167 AMS). The work of the second author is supported by Grant PID2022-137612NB-I00 funded by MCIN/AEI/ 10.13039/501100011033 and ERDF “A way of making Europe.” The third author gratefully acknowledges the support of the NSF (grant DMS-2200850) and the Joshua Barlaz Chair in Mathematics.
1. Introduction
In [ILNT19], Isaacs et al. discovered an unexpected connection between the irrationalities of an irreducible character of a finite group and the parity of its degree: if $\chi$ is an irreducible character of odd degree and its field of values is ${\mathbb{Q}}(\sqrt{d})$ for some square-free integer $d\neq-1$, then $d\equiv 1\pmod{4}$. (Here, the field of values of $\chi$, denoted ${\mathbb{Q}}(\chi)$, is the smallest extension of $\mathbb{Q}$} containing all values of $\chi$.) This result was later generalized to arbitrary higher irrationalities by the second and third authors in [NT21], providing a complete classification of the fields of values of the irreducible characters of odd degree. For odd primes $p$, this description is conjectured, remains open, and is reduced in [NT21] to a problem on quasi-simple groups.
Needless to say, the focus so far has been on characters of degree not divisible by $p$ (or on height zero characters [NRTV24], if one introduces Brauer $p$-blocks), largely motivated by the Galois–McKay conjecture ([N04]) and its possible extensions. An apparently innocent question raised in [IN24] (Question D), however, has shifted this perspective: could the results of [NT21] also be extended to characters of degree divisible by $p$? Given a prime $p$, what are the field of values of the characters divisible by a given power of $p$? The most natural setting to look for examples and patterns is when $p=2$ in quadratic fields ${\mathbb{Q}}(\sqrt{d})$, where $d$ is a square-free integer. In all the cases that we have checked, we observe, for instance, that if ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\sqrt{d})$ with even $d\neq\pm 2$, then $\chi(1)$ is divisible by $4$.
Our first result solves one direction of the classification of the abelian number fields that can occur as fields of values of irreducible characters whose degree has a given $p$-part. Throughout this paper, let $\nu_{p}$ denote the $p$-adic valuation function on the integers; that is, $\nu_{p}(n)=a$ if $p^{a}$ is the largest power of $p$ dividing $n$. If $\mathbb{F}/{\mathbb{Q}}$ is an abelian extension, recall that the conductor $c(\mathbb{F})$ is the smallest integer $n\geq 1$ such that $\mathbb{F}$ is contained in the $n$-th cyclotomic field ${\mathbb{Q}}_{n}:={\mathbb{Q}}(e^{2\pi i/n})$. Also, if $m$ is any positive integer, we let $\mathbb{F}_{m}:=\mathbb{F}(e^{2\pi i/m})$. For a finite group $G$ and $\chi\in{\rm Irr}(G)$, we define $c(\chi)$ to be the conductor of the field ${\mathbb{Q}}(\chi)$.
THEOREM A.
Let $p$ be a prime, $\mathbb{F}$ an abelian extension of $\mathbb{Q}$}, and $a:=\nu_{p}(c(\mathbb{F}))$. Then, for every integer $b\geq\nu_{p}([\mathbb{F}_{p^{a}}:\mathbb{F}])$, there exist a finite group $G$ and $\chi\in\operatorname{Irr}(G)$ such that ${\mathbb{Q}}(\chi)=\mathbb{F}$ and $\nu_{p}(\chi(1))=b$.
To complete the classification, we would need to prove the following, which extends the still unproved Conjecture B.3 of [NT21] on characters of degree not divisible by $p$ to arbitrary characters.
CONJECTURE B.
Let $\chi$ be an irreducible character of some finite group such that $a:=\nu_{p}(c(\chi))\geq 1$. Then $\chi(1)$ is divisible by ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$.
A few remarks are in order. First, the degree of the extension ${\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi)$ can be any power of $p$ when $a$ varies. Indeed: on one hand, every abelian extension $\mathbb{F}/{\mathbb{Q}}$ can arise as a character field (see Lemma 2.3). On the other hand, given a prime power $p^{a}$, elementary number theory shows that there exists an extension $\mathbb{F}\supseteq{\mathbb{Q}}_{p}$ such that $\nu_{p}(c(\mathbb{F}))=a$ and $\mathbb{F}_{p^{a}}:\mathbb{F}]=p^{a-1}$. Second, if ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\sqrt{d})$ with even $d\neq\pm 2$, then $c(\chi)=4|d|$. Hence $\nu_{2}(c(\chi))=3$ and ${\mathbb{Q}}_{2^{3}}(\chi):{\mathbb{Q}}(\chi)]=4$. This would explain our earlier observation on the quadratic character fields and the divisibility by $4$ of character degrees.
Since ${\mathbb{Q}}_{p}(\chi):{\mathbb{Q}}(\chi)]=[{\mathbb{Q}}_{p}:({\mathbb{Q}}_{p}\cap{\mathbb{Q}}(\chi))]$, which divides ${\mathbb{Q}}_{p}:{\mathbb{Q}}]=p-1$, it follows that ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$ is precisely the $p$-part of ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}(\chi)]$. Theorem A therefore implies the following.
THEOREM C.
Let $\mathbb{F}$ be an abelian extension of $\mathbb{Q}$} and $a:=\nu_{p}(c(\mathbb{F}))$. Let $b\in{\mathbb{Z}}^{\geq 0}$. Assume that Conjecture B holds true. Then $\mathbb{F}$ is the field of values of an irreducible character $\chi$ of some finite group with $\nu_{p}(\chi(1))=b$ if and only if
$\nu_{p}([\mathbb{F}_{p^{a}}:\mathbb{F}])\leq b.$
The next result presents evidence in support of Conjecture B.
THEOREM D.
Let $\chi$ be an irreducible character of some finite group $G$. Suppose $a:=\nu_{p}(c(\chi))\geq 1$. Then $\chi(1)$ is divisible by ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$ if one the following cases occurs.
  1. (a)
    $G$ is $p$-solvable;
  2. (b)
    $G$ is an alternating group;
  3. (c)
    $G$ is a general linear or general unitary group.
  4. (d)
    $G$ is the special linear group $\operatorname{SL}_{n}(q)$ and $p\nmid\gcd(n,q-1)$.
  5. (e)
    $G$ is the special unitary group $\operatorname{SU}_{n}(q)$ and $p\nmid\gcd(n,q+1)$.
Also see Theorem 3.9 for results concerning $p^{\prime}$-extensions of $\operatorname{SL}_{n}(q)$ and $\operatorname{SU}_{n}(q)$.
Next we turn to the related problem of classifying the fields of values of the irreducible characters of finite groups whose order has a given $p$-part (in particular, groups of $p^{\prime}$-order). This leads us to the following basic problem: is it always true that $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$? Interestingly, the recently proposed inductive Feit condition ([BKNT25]) implies what seemed to be an intractable problem.
THEOREM E.
Suppose that the finite simple groups satisfy the inductive Feit condition. Then the following hold:
  1. (a)
    If $G$ is a finite group and $\chi\in{\rm Irr}(G)$, then $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$.
  2. (b)
    Let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension, let $p$ be a prime, and $b\in{\mathbb{Z}}_{\geq 0}$. Then $\mathbb{F}={\mathbb{Q}}(\chi)$ for some $\chi\in{\rm Irr}(G)$, where $G$ is a finite group with $\nu_{p}(|G|)=b$, if and only if
    $\nu_{p}(c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b.$
Theorems A, D, and E are proved in Sections 2, 3, and 4, respectively.
2. THEOREM A
Our notation is mostly standard and essentially follows [I76]. We begin by recalling a result of G. M. Cram that will be used frequently.
Lemma 2.1.
Let $G$ be a finite group and $N\trianglelefteq\,G$. Let $\chi\in\operatorname{Irr}(G)$ and $\theta$ be an irreducible constituent of $\chi_{N}$. Let $\psi$ be the Clifford correspondent of $\chi$ with respect to $\theta$. Then ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$ and ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$ divides $\chi(1)/\psi(1)$.
Proof.
This follows from [C88, Lemma 1.1]. ∎
The next lemma easily implies, using direct products, that if $\mathbb{F}$ is the field of values of an irreducible character $\chi$ with $\chi(1)_{p}=p^{b}$ for $b\in{\mathbb{Z}}_{\geq 0}$, then $\mathbb{F}$ also arises as the field of values of a character whose degree has any $p$-part at least $p^{b}$. Therefore, to classify the fields of values of irreducible characters whose degree has a prescribed $p$-part, the question we really have to solve is: given an abelian number field $\mathbb{F}$, what is the smallest possible $p$-part of the degree of an irreducible character $\chi$ such that ${\mathbb{Q}}(\chi)=\mathbb{F}$?
Lemma 2.2.
Let $d\in{\mathbb{Z}}_{\geq 1}$. Then there exists a solvable group $G$ and a rational-valued character $\chi\in\operatorname{Irr}(G)$ such that $\chi(1)=d$.
Proof.
This is [HT23, Lemma 2.1]. ∎
We also need the following.
Lemma 2.3.
Let $n\in{\mathbb{Z}}_{\geq 1}$ and $\mathbb{F}\supseteq{\mathbb{Q}}$ be a subfield of ${\mathbb{Q}}_{n}$. Then there exists an irreducible character $\chi$ (of some finite solvable group) such that ${\mathbb{Q}}(\chi)=\mathbb{F}$ and $\chi(1)=[{\mathbb{Q}}_{n}:\mathbb{F}]$.
Proof.
This is [FG72, Theorem 2.2]. See also [NT21, Theorem 2.2]. ∎
For solvable groups, Theorem C follows from known results.
Theorem 2.4.
Let $\mathbb{F}$ be an abelian extension of $\mathbb{Q}$} and $b\in{\mathbb{Z}}_{\geq 0}$. Then $\mathbb{F}$ is the field of values of an irreducible character $\chi$ of a finite solvable group with $\nu_{p}(\chi(1))=b$ if and only if
$\nu_{p}([{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b.$
Proof.
The ‘only if’ implication directly follows from Cram’s theorem ([C88, Theorem 0.1]). For the reverse implication, let $a:=\nu_{p}([{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b$. By Lemma 2.3, there exists a solvable group $H$ and a character $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\psi)=\mathbb{F}$ and $\psi(1)=[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$. On the other hand, by Lemma 2.2, there exists a solvable group $K$ and a rational-valued character $\varphi\in\operatorname{Irr}(K)$ such that $\varphi(1)=p^{b-a}$.
Let $G:=H\times K$ and $\chi:=\psi\otimes\varphi\in\operatorname{Irr}(G)$. We have ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\psi,\varphi)={\mathbb{Q}}(\psi)=\mathbb{F}$ and $\nu_{p}(\chi(1))=\nu_{p}(\psi(1)\varphi(1))=\nu_{p}(\psi(1))+\nu_{p}(\varphi(1))=a+(b-a)=b$. ∎
The next result we require is the $p^{\prime}$-degree case of Theorem A.
Theorem 2.5 ([NT21], Theorem B2).
Let $\mathbb{F}$ be an abelian extension of $\mathbb{Q}$} with $\nu_{p}(c(\mathbb{F}))=a$. Suppose that ${\mathbb{Q}}_{p^{a}}:(\mathbb{F}\cap{\mathbb{Q}}_{p^{a}})]$ is not divisible by $p$. Then there exists a finite group $G$ and $\chi\in\operatorname{Irr}_{p^{\prime}}(G)$ such that ${\mathbb{Q}}(\chi)=\mathbb{F}$.
One of the key constructions in producing the desired group and character in Theorem 2.5 is to reduce from a character field of values to a smaller one where the extension is cyclic, see [NT21, Theorem 3.3]. For our purpose, we need to generalize this to arbitrary (abelian) extensions.
Theorem 2.6.
Let $\chi$ be an irreducible character of a finite group $G$ and set ${\mathbb{E}}:={\mathbb{Q}}(\chi)$. Then, for any subfield $\mathbb{F}$ of $\mathbb{E}$}, there exists a finite group $H$ and a character $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\psi)=\mathbb{F}$ and $\psi(1)=n\chi(1)^{n}$, where $n:=[{\mathbb{E}}:\mathbb{F}]$.
Proof.
Write $\mathcal{G}:=\mathrm{Gal}({\mathbb{E}}/\mathbb{F})$ as a direct product ${\sf C}_{n_{1}}\times{\sf C}_{n_{2}}\times\cdots\times{\sf C}_{n_{k}}$ of cyclic groups of orders $n_{1},...,n_{k}$ with $n_{i}>1$ and $n_{i}$ is divisible by $n_{i+1}$. Note that $n:=\prod_{i=1}^{k}n_{i}=|\mathcal{G}|$. Let $\tau_{i}$ be a generator of ${\sf C}_{i}$, so that
$\mathcal{G}=\{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}\mid 1\leq a_{i}\leq n_{i}\text{ for every }1\leq i\leq k\}$
Consider the $n$-iterated direct product of $G$ and view its elements as $k$-dimensional matrices:
$G^{n}:=\{X=(x_{a_{1},...,a_{k}})_{1\leq a_{i}\leq n_{i}\text{ for every }1\leq i\leq k}\mid x_{a_{1},...,a_{k}}\in G\}.$
Let $\tau_{i}$ act on $G^{n}$ by permuting the $i$th-dimensional indices of the entries $x_{a_{1},...,a_{k}}$ but fixing the indices at the other dimensions, in the following way
$(x_{a_{1},...,a_{k}})^{\tau_{i}}=(x_{a_{1},...,a_{i-1},a_{i}-1,a_{i+1},...,a_{k}}),$
where two indices at the $i$th-dimension are identified if they have the same residue modulo $n_{i}$. This action induces an embedding of $\mathcal{G}={\sf C}_{n_{1}}\times{\sf C}_{n_{2}}\times\cdots\times{\sf C}_{n_{k}}$ into the symmetric group ${\mathrm{Sym}}_{n}$. Let
$H:=G^{n}\rtimes\mathcal{G}$
be the corresponding semidirect product.
Let
$\varphi:=\bigotimes_{1\leq a_{i}\leq n_{i}}\varphi_{a_{1},...,a_{k}}\in\operatorname{Irr}(G^{n}),$
where
$\varphi_{a_{1},...,a_{k}}=\chi^{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}}\in\operatorname{Irr}(G);$
that is,
$\varphi(X)=\prod_{1\leq a_{i}\leq n_{i}\atop 1\leq i\leq k}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...,a_{k}}).$
The action of a typical element $\tau=\tau_{1}^{b_{1}}\cdots\tau_{k}^{b_{k}}$ of $\mathcal{G}$ on $\varphi$ is given by
$(\varphi^{\tau})_{a_{1},...,a_{k}}=\chi^{\tau_{1}^{a_{1}+b_{1}}\cdots\tau_{k}^{a_{k}+b_{k}}}.$
It is clear that $\varphi^{\tau}\neq\varphi$ whenever $\tau\neq 1$. Therefore, the stabilizer $I_{H}(\varphi)$ of $\varphi$ in $H$ is simply $G^{n}$, which in turn implies that
$\psi:=\varphi^{H}\in\operatorname{Irr}(H).$
We have $\psi(1)=[H:G^{n}]\varphi(1)=n\chi(1)^{n}$. The rest of the proof will establish that ${\mathbb{Q}}(\psi)=\mathbb{F}$. This can be done by an suitable modification of the arguments in the proof of [NT21, Theorem 3.3].
By the definition of $\varphi$, we know ${\mathbb{Q}}(\varphi)={\mathbb{Q}}(\chi)={\mathbb{E}}$. The character-induction formula then implies that ${\mathbb{Q}}(\psi)\subseteq{\mathbb{E}}$. Furthermore, for each $X\in G^{n}$ and $j\in\{1,...,k\}$,
$\varphi(X^{\tau_{j}})=\prod_{1\leq a_{i}\leq n_{i}\atop 1\leq i\leq k}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...a_{j-1},a_{j}-1,a_{j+1},...,a_{k}}).$
On the other hand,
$\varphi^{\tau_{j}}(X)=\prod_{1\leq a_{i}\leq n_{i}\atop 1\leq i\leq k}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{j-1}^{a_{j-1}}\tau_{j}^{a_{j}+1}\tau_{j+1}^{a_{j+1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...,a_{k}}).$
If all the $a_{1},...,a_{j-1},a_{j+1},...,a_{k}$ are fixed, the sub-products along the $j$th dimension are equal:
$\prod_{1\leq a_{j}\leq n_{j}}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...a_{j-1},a_{j}-1,a_{j+1},...,a_{k}})=\prod_{1\leq a_{j}\leq n_{j}}\chi^{\tau_{1}^{a_{1}}\cdots\tau_{j-1}^{a_{j-1}}\tau_{j}^{a_{j}+1}\tau_{j+1}^{a_{j+1}}\cdots\tau_{k}^{a_{k}}}(x_{a_{1},...,a_{k}})$
It follows that
$\varphi(X^{\tau_{j}})=\varphi^{\tau_{j}}(X).$
Since $\mathcal{G}$ is generated by $\tau_{j},1\leq j\leq k$, we conclude that
$\varphi(X^{\tau})=\varphi^{\tau}(X)$
for every $\tau\in\mathcal{G}$. That is, the Galois action of $\mathcal{G}$ on $\varphi$ is compatible with our defined action of $\mathcal{G}$ on $G^{n}$.
Now, for every $X\in G^{n}$ and $\tau\in\mathcal{G}$, by the Clifford theorem, we have
$\psi(X)^{\tau}=$ $\left(\sum_{\tau^{\prime}\in\mathcal{G}}\varphi(X^{\tau^{\prime}})\right)^{\tau}=\left(\sum_{\tau^{\prime}\in\mathcal{G}}\varphi(X)^{\tau^{\prime}}\right)^{\tau}$
$=$ $\sum_{\tau^{\prime}\in\mathcal{G}}\varphi(X)^{\tau\tau^{\prime}}=\sum_{\tau^{\prime}\in\mathcal{G}}\varphi(X)^{\tau^{\prime}}=\psi(X).$
Therefore, $\psi(X)$ is fixed by $\mathcal{G}$ point-wisely. As $\psi$ vanishes outside $G^{n}$, we deduce that ${\mathbb{Q}}(\psi)={\mathbb{Q}}(\psi_{G^{n}})$ is contained in the fixed field of $\mathcal{G}$ in $\mathbb{E}$}. That is, ${\mathbb{Q}}(\psi)\subseteq\mathbb{F}$.
By Lemma 2.1, we have ${\mathbb{Q}}(\varphi):{\mathbb{Q}}(\psi)]$ divides $H:G^{n}]=n$. As mentioned already that ${\mathbb{Q}}(\varphi)={\mathbb{E}}$, it follows that ${\mathbb{E}}:{\mathbb{Q}}(\psi)]\leq n=[{\mathbb{E}}:\mathbb{F}]$. This and the conclusion of the preceding paragraph imply that ${\mathbb{Q}}(\psi)=\mathbb{F}$. The proof is complete. ∎
We can now complete the proof of Theorem A, which is restated.
Theorem 2.7.
Let $\mathbb{F}$ be an abelian extension of $\mathbb{Q}$} and $a:=\nu_{p}(c(\mathbb{F}))$. Let $c:=\nu_{p}([\mathbb{F}_{p^{a}}:\mathbb{F}])$. Then, for every integer $b\geq c$, there exists a finite group $G$ and a character $\chi\in\operatorname{Irr}(G)$ such that ${\mathbb{Q}}(\chi)=\mathbb{F}$ and $\nu_{p}(\chi(1))=b$.
Proof.
Let ${\mathbb{E}}:={\mathbb{Q}}_{p^{a}}(\mathbb{F})$. It is clear that $\nu_{p}(c({\mathbb{E}}))=a$ and ${\mathbb{Q}}_{p^{a}}:({\mathbb{E}}\cap{\mathbb{Q}}_{p^{a}})]=1$. Using Theorem 2.5, we obtain a finite group $H$ and an irreducible character $\psi$ of $H$ of $p^{\prime}$-degree such that ${\mathbb{Q}}(\psi)={\mathbb{E}}$.
By the hypothesis, we have $\mathbb{F}_{p^{a}}:\mathbb{F}]=[{\mathbb{E}}:\mathbb{F}]=p^{c}f$ for some $f\in{\mathbb{Z}}^{+}$ not divisible by $p$. Now applying Theorem 2.6 to the extension ${\mathbb{E}}/\mathbb{F}$, we then obtain a finite group $M$ and $\varphi\in\operatorname{Irr}(M)$ such that
$\varphi(1)=p^{c}f\psi(1)^{p^{c}f}\text{ and }{\mathbb{Q}}(\varphi)=\mathbb{F}.$
Next, we use Lemma 2.2 to have a finite group $N$ and some rational-valued character $\lambda\in\operatorname{Irr}(N)$ such that $\lambda(1)=p^{b-c}$.
Finally, let $G:=M\times N$ and $\chi:=\varphi\otimes\lambda$ – an irreducible character of $G$. This $\chi$ is indeed the required character. To see that, we have
$\chi(1)=\varphi(1)\lambda(1)=p^{c}f\psi(1)^{p^{c}f}\cdot p^{b-c}=p^{b}f\psi(1)^{p^{c}f}.$
Recall that both $f$ and $\psi(1)$ are not divisible by $p$. Therefore $\nu_{p}(\chi(1))=b$. Moreover,
${\mathbb{Q}}(\chi)={\mathbb{Q}}(\varphi,\lambda)={\mathbb{Q}}(\varphi)=\mathbb{F},$
as desired. ∎
3. Theorem D
We begin with the following lemma about $p$-groups.
Lemma 3.1.
Suppose that $P$ is a $p$-group. Let $\alpha\in{\rm Irr}(P)$ be with $c(\alpha)=p^{a}$ and $a\geq 1$.
  1. (a)
    If $p$ is odd or $p^{a}=4$, then ${\mathbb{Q}}_{p^{a}}(\alpha)={\mathbb{Q}}_{p}(\alpha)$.
  2. (b)
    If $p=2$ and $p^{a}\geq 8$, then $|{\mathbb{Q}}_{p^{a}}(\alpha):{\mathbb{Q}}_{p}(\alpha)|\leq 2$ divides $\alpha(1)$.
Proof.
If $p$ is odd, then we know that ${\mathbb{Q}}(\alpha)={\mathbb{Q}}_{p^{a}}$ by [NT21, Theorem 2.3]. If $p=2$ and $a=2$, then ${\mathbb{Q}}(\alpha)\subseteq{\mathbb{Q}}(i)$ and necessarily ${\mathbb{Q}}(\alpha)={\mathbb{Q}}(i)$, so there is nothing to prove. Assume now that $p=2$ and $a\geq 3$. By elementary Galois theory, we know that ${\mathbb{Q}}_{2^{a}}/{\mathbb{Q}}_{8}$ is a cyclic extension with intermediate fields ${\mathbb{Q}}_{8}\subseteq{\mathbb{Q}}_{16}\subseteq\ldots\subseteq{\mathbb{Q}}_{2^{a}}$. Therefore ${\mathbb{Q}}_{8}(\alpha)={\mathbb{Q}}_{2^{a}}={\mathbb{Q}}_{2^{a}}(\alpha)$. Also ${\mathbb{Q}}_{8}\cap{\mathbb{Q}}(\alpha)$ cannot be $\mathbb{Q}$}, since ${\mathbb{Q}}_{2^{a}}$ has only three subfields of degree 2 and all of them are contained in ${\mathbb{Q}}_{8}$. If $\alpha$ is linear, notice that ${\mathbb{Q}}_{p^{a}}(\alpha):{\mathbb{Q}}_{p}(\alpha)]=1$. Part (b) easily follows. ∎
Theorem 3.2.
Let $G$ be a $p$-solvable finite group, and let $\chi\in{\rm Irr}(G)$ be with $a=\nu_{p}(c(\chi))\geq 1$. Then ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$ divides $\chi(1)$.
Proof.
Notice that ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]$ is a power of $p$. We argue by induction on $|G|$. Suppose that there is $N\trianglelefteq\,G$, and $\eta^{G}=\chi$, where $\eta\in{\rm Irr}(V)$ is a Clifford correspondent of $\chi$. By Lemma 2.1, we know that ${\mathbb{Q}}(\eta):{\mathbb{Q}}(\chi)]$ divides $|G:V|$. We have that $c(\chi)$ divides $c(\eta)$. Write $c(\eta)=p^{b}$, where $b\geq 1$. By induction, we have that ${\mathbb{Q}}_{p^{a}}(\eta):{\mathbb{Q}}_{p}(\eta)]$ divides ${\mathbb{Q}}_{p^{b}}(\eta):{\mathbb{Q}}_{p}(\eta)]$ divides $\eta(1)$. If $K={\mathbb{Q}}_{p^{a}}(\chi)\cap{\mathbb{Q}}_{p}(\eta)$, by elementary Galois theory, we have that $K:{\mathbb{Q}}_{p}(\chi)]$ divides ${\mathbb{Q}}(\eta):{\mathbb{Q}}(\chi)]$ which divides $|G:V|$, and we are done.
We may now assume that $\chi$ is quasi-primitive. Hence, $\chi=\alpha\beta$, where $\alpha$ is $p$-special and $\beta$ is $p^{\prime}$-special (see [I18, Theorem 2.9]). We claim that ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]=1$, in this case. First, we prove that ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\alpha,\beta)$. Certainly, ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\alpha,\beta)$. However, if $\sigma\in{\rm Gal}({\mathbb{Q}}(\alpha,\beta)/{\mathbb{Q}}(\chi)$, then we obtain that $\alpha^{\sigma}=\alpha$ and $\beta^{\sigma}=\beta$, by the uniqueness of the product decomposition. Since ${\mathbb{Q}}(\alpha)={\mathbb{Q}}(\alpha_{P})$, where $P\in{\rm Syl}_{p}(G)$ (by the uniqueness of the restriction from $p$-special characters), we have that $c(\alpha\beta)=c(\alpha)c(\beta)$. If $p$ is odd or $p^{a}=4$, then ${\mathbb{Q}}_{p^{a}}={\mathbb{Q}}_{p}(\alpha)$, by Lemma 3.1(a). Then ${\mathbb{Q}}_{p^{a}}(\chi)={\mathbb{Q}}_{p^{a}}(\alpha,\beta)={\mathbb{Q}}_{p^{a}}(\beta)$ and ${\mathbb{Q}}_{p}(\chi)={\mathbb{Q}}_{p}(\alpha,\beta)={\mathbb{Q}}_{p^{a}}(\beta)$. Suppose finally that $p=2$ and $a\geq 3$. If $\chi$ has odd-degree, then the theorem follows from [NT21, Theorem A1]. So we may assume that $\chi(1)$ is even. By Lemma 3.1, we have that ${\mathbb{Q}}_{p^{a}}(\chi):{\mathbb{Q}}_{p}(\chi)]\leq 2$, and the theorem follows. ∎
Theorem 3.3.
Let $n\geq 1$ and $\chi\in{\rm Irr}({\mathrm{Alt}}_{n})\cup{\rm Irr}({\mathrm{Sym}}_{n})$. Then $\nu_{p}(c(\chi))\leq 1$. In particular, Theorem D holds for the symmetric and alternating groups.
Proof.
It is well known that the characters of ${\mathrm{Sym}}_{n}$ are all rational, so the result is immediate in this case. Hence, we may assume that $\chi\in\operatorname{Irr}({\mathrm{Alt}}_{n})$. For background on the ordinary irreducible characters of ${\mathrm{Alt}}_{n}$, we refer the reader to [JK81, §5].
Suppose that $\Psi$ is the irreducible character of ${\mathrm{Sym}}_{n}$ lying over $\chi$, and let $\lambda$ be the partition of $n$ labeling $\Psi$. If $\lambda$ is not self-conjugate, then $\chi=\Psi_{{\mathrm{Alt}}_{n}}$ and the result is again trivial. Thus we assume that $\lambda$ is self-conjugate. In this case, $\Psi_{{\mathrm{Alt}}_{n}}=\chi+\chi^{\sigma},$ where $\sigma$ is an odd permutation in ${\mathrm{Sym}}_{n}$.
Let $h(\lambda)=(h_{11}^{\lambda},\dots,h_{kk}^{\lambda})$ denote the partition of $n$ whose parts $h_{ii}^{\lambda}$ are the hook lengths at the diagonal nodes $(i,i)$ of the Young diagram of $\lambda$. Here $k$ is the length of the main diagonal. The conjugacy class of ${\mathrm{Sym}}_{n}$ of cycle type $h(\lambda)$ splits into two conjugacy classes of ${\mathrm{Alt}}_{n}$, say $h(\lambda)^{+}$ and $h(\lambda)^{-}$.
According to [JK81, Theorem 2.5.13], the values of $\chi$ (and of $\chi^{\sigma}$) are rational at all elements, except possibly at those belonging to $h(\lambda)^{\pm}$. In fact, one has
${\mathbb{Q}}(\chi)={\mathbb{Q}}\!\left(\sqrt{(-1)^{(n-k)/2}\,h_{11}^{\lambda}\cdots h_{kk}^{\lambda}}\right).$
Let $d$ denote the square-free part of $h_{11}^{\lambda}\cdots h_{kk}^{\lambda}$, that is, the product of the prime divisors occurring with odd exponents. Note that all the hook lengths $h_{ii}^{\lambda}$ are odd and that
$n-k=\sum_{i=1}^{k}(h_{ii}^{\lambda}-1)$
is even. We shall use $c(z)$ to denote the conductor of a sum of roots of unity $z$, which is the smallest positive integer $t$ such that $z\in{\mathbb{Q}}_{t}$. If $n-k\equiv 0\pmod{4}$, then $h_{11}^{\lambda}\cdots h_{kk}^{\lambda}\equiv 1\pmod{4}$, and hence
$c(\chi)=c\!\left(\sqrt{h_{11}^{\lambda}\cdots h_{kk}^{\lambda}}\right)=c(\sqrt{d})=d.$
On the other hand, if $n-k\equiv 2\pmod{4}$, then $h_{11}^{\lambda}\cdots h_{kk}^{\lambda}\equiv 3\pmod{4}$, and hence
$c(\chi)=c\!\left(\sqrt{-h_{11}^{\lambda}\cdots h_{kk}^{\lambda}}\right)=c(\sqrt{-d})=d.$
It follows that $\nu_{p}(c(\chi))=1$ if $p$ divides $d$, and $\nu_{p}(c(\chi))=0$ otherwise. This completes the proof. ∎
We now prove Theorem D for finite general linear groups $G=\operatorname{GL}_{n}(q)$, where $n\geq 1$ and $q$ is a prime power, with a natural module $V=\mathbb{F}_{q}^{n}=\langle e_{1},\ldots,e_{n}\rangle_{\mathbb{F}_{q}}$. It is convenient for us to use the Dipper-James classification of complex irreducible characters of $G$, as described in [J86]. Namely, every $\chi\in\operatorname{Irr}(G)$ can be written uniquely, up to a permutation of the pairs $\{(s_{1},\lambda_{1}),\dots,(s_{m},\lambda_{m})\}$, in the form
(3.1) $\chi=S(s_{1},\lambda_{1})\circ S(s_{2},\lambda_{2})\circ\ldots\circ S(s_{m},\lambda_{m}).$
Here, $s_{i}\in\overline{\mathbb{F}_{q}}^{\times}$ has degree $d_{i}$ over $\mathbb{F}_{q}$, $\lambda_{i}\vdash k_{i}$, $\sum^{m}_{i=1}k_{i}d_{i}=n$, and the $m$ elements $s_{i}$ have pairwise distinct minimal polynomials over $\mathbb{F}_{q}$. In particular, $S(s_{i},\lambda_{i})$ is an irreducible character of $\operatorname{GL}_{k_{i}d_{i}}(q)$. Furthermore, there is a parabolic subgroup $P_{\chi}=U_{\chi}L_{\chi}$ of $G$ with Levi subgroup $L_{\chi}=\operatorname{GL}_{k_{1}d_{1}}(q)\times\ldots\times\operatorname{GL}_{k_{m}d_{m}}(q)$ and unipotent radical $U_{\chi}$. The (outer) tensor product
$\psi:=S(s_{1},\lambda_{1})\otimes S(s_{2},\lambda_{2})\otimes\ldots\otimes S(s_{m},\lambda_{m})$
is an $L_{\chi}$-character, and $\chi$ is obtained from $\psi$ via the Harish-Chandra induction $R^{G}_{L_{\chi}}$, i.e. we first inflate $\psi$ to a $P_{\chi}$-character and then induce it to $G$.
Lemma 3.4.
In the above notation, assume that the integer $N\geq 1$ is divisible by the order of every $s_{i}$, $1\leq i\leq m$, and that $\sigma\in\mathrm{Gal}({\mathbb{Q}}_{N}/{\mathbb{Q}})$ sends a primitive $N^{\mathrm{th}}$ root of unity $\zeta\in{\mathbb{C}}^{\times}$ to $\zeta^{k}$, $1\leq k<N$. Then for the character $\chi$ in (3.1),
$\sigma(\chi)=S(s_{1}^{k},\lambda_{1})\circ S(s_{2}^{k},\lambda_{2})\circ\ldots\circ S(s_{m}^{k},\lambda_{m}).$
Proof.
First we show that $s_{i}^{k}$ also has degree $d_{i}$ over $\mathbb{F}_{q}$, for any $i$. Indeed, $s_{i}\in\mathbb{F}_{q^{d_{i}}}$ implies that $s_{i}^{k}\in\mathbb{F}_{q^{d_{i}}}$, and hence $s_{i}^{k}$ has degree $d^{\prime}|d_{i}$. Conversely, as $\gcd(k,N)=1$, we can find $l\in{\mathbb{Z}}_{\geq 1}$ such that N|(kl1)N|(kl-1). Now $s_{i}=s_{i}^{kl}\in\mathbb{F}_{q^{d^{\prime}}}$, and so $d_{1}|d^{\prime}$, whence $d^{\prime}=d_{i}$.
Since the Harish-Chandra induction $R^{G}_{L_{\chi}}$ commutes with Galois action, it suffices to prove the claim in the case $m=1$. Now, if $d_{1}>1$, then $\chi=S(s_{1},\lambda_{1})$ is Lusztig induced from the character $S(s_{1},\lambda_{1})$ of the Levi subgroup $L=\operatorname{GL}_{k_{1}}(q^{d_{1}})$ of $G=\operatorname{GL}_{k_{1}d_{1}}(q)$, see [FS82, p. 116]. Since the Lefschetz numbers are all integers (see e.g. [DM91, p. 140]), Green functions involved in the Lusztig induction $R^{G}_{L}$ are rational-valued, and thus $R^{G}_{L}$ also commutes with Galois actions. Thus it suffices to prove the claim in the case $m=d_{1}=1$. In the latter case, $\chi=S(1,\lambda_{1})\hat{s_{1}}$, where $S(1,\lambda_{1})$ is a rational (unipotent) character, and $\hat{s_{1}}$ is a linear character of $G/[G,G]$ of order equal to the order of $s_{1}$, see [FS82, (1.16)], so we are done. ∎
Lemma 3.5.
Let $p$ be a prime and $A={\sf C}_{p^{n}}\times{\sf C}_{p}$ for some $n\geq 1$. If $B$ is a subgroup of $A$, then there is some $1\leq m\leq n$ such that $B\cong\mathsf{C}_{p^{m}}$ or $B\cong\mathsf{C}_{p^{m}}\times\mathsf{C}_{p}$.
Proof.
Write $A=A_{1}\times A_{2}$ with $A_{1}=\langle x\rangle\cong\mathsf{C}_{p^{n}}$ and $A_{2}=\langle y\rangle\cong\mathsf{C}_{p}$, and consider the natural projection $\pi:A\to A_{1}$ with kernel $A_{2}$. If $\pi|_{B}$ is injective then $B$ is cyclic of order $p^{m}$ with $m\leq n$, and we are done. Otherwise $1\neq\operatorname{Ker}(\pi|_{B})$, so $y\in B$, whence $B=A_{2}\times(A_{1}\cap B)$, and we are done again. ∎
Theorem 3.6.
Theorem D holds for finite general linear groups $G=\operatorname{GL}_{n}(q)$ and finite general unitary groups $\operatorname{GU}_{n}(q)$.
Proof.
Since the statement is trivial when $a=1$ or $n=1$, we may assume that $a,n\geq 2$.
The case of $\operatorname{GU}_{n}(q)$ is entirely similar to that of $\operatorname{GL}_{n}(q)$, changing only $q$ to q-q (in character degrees and subgroup orders, as well as in the definition of $[t]$ for $t\in\overline{\mathbb{F}_{q}}^{\times}$) and Harish-Chandra induction to Lusztig induction. So we will provide the details only for $G=\operatorname{GL}_{n}(q)$. Note that $p\nmid q$ since (3.1) shows that $c(\chi)$ divides $N:=\mathrm{lcm}(o(s_{1}),\ldots,o(s_{m}))$ which divides $\prod^{n}_{i=1}(q^{i}-1)$. So $q$ is power of some prime $r\neq p$.
(a) Here we show that if a $p$-element $\sigma\in\mathrm{Gal}({\mathbb{Q}}_{N}/{\mathbb{Q}}_{p}(\chi))$ induces an element of $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$ of order $p^{a_{1}}$ for some $a_{1}\geq 0$, then $\chi(1)$ is divisible by $p^{a_{1}}$, in fact by $2^{a_{1}+1}$ if $p=2\leq a_{1}$.
The case $a_{1}=0$ is trivial. so we may assume $a_{1}\geq 1$, and note that $\chi$ is $\sigma$-invariant. Writing $\chi$ in the form (3.1) and applying Lemma 3.4 with $\sigma(\zeta)=\zeta^{f}$, we see that the induced map $\tau:S(s,\lambda)\mapsto S(s^{f},\lambda)$ acts on the set $S$ of the characters $S(s_{i},\lambda_{i})$, $1\leq i\leq m$.
(a1) First we show that if some $\tau$-orbit $\mathcal{O}$} on $S$ has length $p^{b}$ with $b\geq 0$, say $S(s_{1},\lambda_{1})\in{\mathcal{O}}$, and $p$ divides $q^{k}-1$ with $k:=k_{1}d_{1}$, then both
$N_{1}:=\frac{|\operatorname{GL}_{kp^{b}}(q)|_{r^{\prime}}}{|\operatorname{GL}_{k}(q)|_{r^{\prime}}^{p^{b}}}$
and $\chi(1)$ are divisible by $p^{b}$, in fact by $p^{b+1}$ if $p=2\leq b$.
Indeed, the statement is obvious if $b=0$. Assume $b\geq 1$. Then for any $1\leq i\leq b$, $p|(q^{kp^{i}}-1)/(q^{kp^{i-1}}-1)$, hence $p^{b}$ divides $(q^{kp^{b}}-1)/(q^{k}-1)$. Since
$N_{1}:=N_{2}\cdot\frac{q^{kp^{b}}-1}{q^{k}-1},\mbox{ where }N_{2}:=\frac{|\operatorname{GL}_{kp^{b}-1}(q)|_{r^{\prime}}}{|\operatorname{GL}_{k}(q)|_{r^{\prime}}^{p^{b}-1}|\operatorname{GL}_{k-1}(q)|_{r^{\prime}}}$
is an integer, $p^{b}|N_{1}$. Next, $N_{1}$ divides
$\frac{|\operatorname{GL}_{n}(q)|_{r^{\prime}}}{\prod^{m}_{i=1}|\operatorname{GL}_{k_{i}d_{i}}(q)|_{r^{\prime}}}.$
So $N_{1}|\chi(1)$, and hence $p^{b}|\chi(1)$.
Assume in addition that $p=2\leq b$. Then we can express $N_{2}$ as
$\biggl(\frac{|\operatorname{GL}_{k\cdot 2^{b-1}-1}(q)|_{r^{\prime}}}{|\operatorname{GL}_{k}(q)|_{r^{\prime}}^{2^{b-1}-1}|\operatorname{GL}_{k-1}(q)|_{r^{\prime}}}\biggr)^{2}\cdot\frac{|\operatorname{GL}_{k\cdot 2^{b}-1}(q)|_{r^{\prime}}}{|\operatorname{GL}_{k\cdot 2^{b-1}}(q)|_{r^{\prime}}|\operatorname{GL}_{k\cdot 2^{b-1}-1}(q)|_{r^{\prime}}}\cdot\frac{q^{k\cdot 2^{b-1}}-1}{q^{k}-1},$
where the first two factors are integers, and the third one is an even integer. So $2^{b+1}$ divides both $N_{1}$ and $\chi(1)$.
(a2) For any $t\in\overline{\mathbb{F}_{q}}^{\times}$ of degree say $l$ over $\mathbb{F}_{q}$, let $[t]$ denote the set $\{t,t^{q},\ldots,t^{q^{l-1}}\}$. Suppose that some $\tau$-orbit $\mathcal{O}$} on $S$ has length $p^{b}$ with $b\geq 0$, say $S(s_{1},\lambda_{1})\in{\mathcal{O}}$, and $p|(q^{d}-1)$ with $d:=d_{1}$. As $\tau^{p^{b}}$ fixes $S(s_{1},\lambda_{1})$, $\tau^{p^{b}}$ acts on the set $s_{1}]$. Suppose in addition that some $\tau^{p^{b}}$-orbit on $s_{1}]$ has length $p^{c}$ for some $c\geq 0$. Then we show that $p^{b+c}|\chi(1)$, and in fact $2^{b+c+1}|\chi(1)$ if $p=2\leq b+c$.
Indeed, write $k:=k_{1}$. If $c=0$, then we are done by (a1). So we assume that $c\geq 1$. By (a1),
$N_{1}:=\frac{|\operatorname{GL}_{kdp^{b}}(q)|_{r^{\prime}}}{|\operatorname{GL}_{kd}(q)|_{r^{\prime}}^{p^{b}}}$
is divisible by $p^{b}$, in fact by $p^{b+1}$ if $p=2\leq b$.
Since the maps $\tau^{p^{b}}:x\mapsto x^{f^{p^{b}}}$ and $x\mapsto x^{q}$ have commuting actions on $s_{1}]$, with the latter acting transitively, we see that $p^{c}$ divides the size of any $\tau^{p^{b}}$-orbit on $s_{1}]$, which has cardinality $d$. Thus $p^{c}|d$; write $d=p^{c}d_{1}$. As $p\nmid q$, we have $q^{d}\equiv q^{d_{1}}\pmod{p}$, and so $p|(q^{d_{1}}-1)$. Hence $p$ divides each $q^{id_{1}}-1$ with $1\leq i\leq p^{c}-1$, and so $p^{p^{c}-1}$ divides $(q-1)(q^{2}-1)\ldots(q^{d-1}-1)$, which is a divisor of
$\frac{(q-1)(q^{2}-1)\ldots(q^{kd}-1)}{(q^{d}-1)(q^{2d}-1)\ldots(q^{kd}-1)}.$
The latter divides the degree of $S(s_{1},\lambda_{1})$, see [KT, Lemma 5.7(ii)]. Also, $c\geq 1$ implies that $p^{c}-1\geq 2^{c}-1\geq c$. It follows that the degree of $S(s_{1},\lambda_{1})$ is divisible by $p^{c}$, in fact by $p^{c+1}$ if $p=2\leq c$.
Thus $p^{b+c}|\chi(1)$; moreover $p^{b+c+1}|\chi(1)$ if $p=2\leq\max(b,c)$. It remains to consider the case $p=2$ and $b=c=1$. In this case $\chi(1)$ is divisible by $N_{1}\bigl(\deg(S(s_{1},\lambda_{1}))\bigr)^{2}$, so by $2^{3}$ as stated.
(a3) Now, if $\tau$ has some orbit of length $p^{b}$ on $S$ containing $S(s_{j},\lambda_{j})$ with $p|o(s_{j})$ and $\tau^{p^{b}}$ has some orbit of length $p^{c}$ on the set $s_{j}]$ with $b+c\geq a_{1}$, then we are done by (a2). So we may assume that $\tau_{1}:=\tau^{p^{a_{1}-1}}$ fixes every $s_{i}$ with $p|o(s_{i})$. Thus if $\tau_{1}(x)=x^{f_{1}}$, then $f_{1}-1$ is divisible by $o(s_{i})$ whenever $p|o(s_{i})$. On the other hand, (3.1) shows that $c(\chi)$ divides $\mathrm{lcm}(o(s_{1}),\ldots,o(s_{m}))$. It follows that $p^{a}=c(\chi)_{p}$ divides $f_{1}-1$. But in this case, $\tau_{1}$ acts trivially on both ${\mathbb{Q}}_{p^{a}}$ and ${\mathbb{Q}}(\chi)$, contrary to the assumption that $\tau_{1}$ has order $p$ in $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$.
(b) By [W09, Theorem 3.4.3], the restriction to ${\mathbb{Q}}_{p^{a}}$ gives an injective homomorphism $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))\to\mathrm{Gal}({\mathbb{Q}}_{p^{a}}/{\mathbb{Q}}_{p})$, and the latter is isomorphic to $\mathsf{C}_{p^{a-1}}$ if $p>2$ and to $\mathsf{C}_{2^{a-2}}\times\mathsf{C}_{2}$ if $p=2$. It follows that $|\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))|=p^{e}$ with $e\leq a-1$.
It suffices to prove the statement in the case $e\geq 1$. Now, if $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$ is cyclic, then we can find an element of order $p^{e}$ in $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$, and so $p^{e}|\chi(1)$ by (a), and we are done.
By Lemma 3.5, it remains to consider the case $p=2\leq e$ and $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))\cong\mathsf{C}_{2^{e-1}}\times\mathsf{C}_{2}$. Now, if $e\geq 3$, then then we can find an element of order $2^{e-1}\geq 4$ in $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$, and so $2^{e+1}|\chi(1)$ by (a), and we are done again. Suppose now that $e=2$ but $4\nmid\chi(1)$. The analysis in (a) then shows that, for every $\tau$ induced by a nontrivial element in $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))$, $\tau$ has exactly one orbit of length $2$ on
$S^{\prime}:=\{(s_{i}^{q^{j}},\lambda_{i})\mid 1\leq i\leq m,0\leq j\leq d_{j}-1\},$
with all other orbits having length $1$. However, since $4\nmid\chi(1)$, the same holds for the induced action of $\mathrm{Gal}({\mathbb{Q}}_{p^{a}}(\chi)/{\mathbb{Q}}_{p}(\chi))=\langle\sigma_{1},\sigma_{2}\rangle\cong\mathsf{C}_{2}^{2}$. Thus if $\tau_{i}$ denotes the induced action of $\sigma_{i}$, $i=1,2$, then we may assume that both $\tau_{1,2}$ act as the same $2$-cycle on $S^{\prime}$. But then the automorphism $\sigma_{1}\sigma_{2}$ of order $2$ acts trivially on $S^{\prime}$, a contradiction. ∎
We will need the following refinement of [NT21, Lemma 4.2]:
Proposition 3.7.
Let $p$ be a prime and let $N$ be a normal subgroup of a finite group $G$ with $p\nmid|G/N|$. Let $\theta\in\operatorname{Irr}(N)$, $J:=\mathrm{Stab}_{G}(\theta)$, and let $\mathbb{F}$ by any extension of $\mathbb{Q}$} inside ${\mathbb{Q}}_{|G|}$. Then the following statements hold.
  1. (i)
    Suppose that $\chi\in\operatorname{Irr}(G|\theta)$, $a=\nu_{p}(c(\chi))$, and $b=\nu_{p}(c(\theta))$. If $a,b\geq 1$ or $p=2$, then $a=b$. If $a=0$ then $b\leq 1$, and if $b=0$ then $a\leq 1$. Furthermore, $p\nmid[\mathbb{F}(\theta):(\mathbb{F}(\chi)\cap\mathbb{F}(\theta))]$. In fact, every prime divisor of $\mathbb{F}(\theta):(\mathbb{F}(\chi_{N})\cap\mathbb{F}(\theta))]$ divides $|G/N|$.
  2. (ii)
    Suppose that all the characters in $\operatorname{Irr}(J|\theta)$ have the same degree (this holds if, for instance, $J/N$ is cyclic). Then, there exists some $\chi\in\operatorname{Irr}(G|\theta)$ such that $p\nmid[\mathbb{F}(\chi):(\mathbb{F}(\chi)\cap\mathbb{F}(\theta))]$.
Proof.
(i) The first statement is just [NT21, Lemma 4.2(ii)]. The second statement follows from [H24, Lemma 2.4(ii)], where it was shown that if $\max\{a,b\}\geq 2$, then $a=b$.
For the third statement, by Clifford’s theorem we have $\chi_{N}=e\sum^{t}_{i=1}\theta_{i}$, where $t=[G:J]$ is coprime to $p$, and $\theta=\theta_{1},\theta_{2},\ldots,\theta_{t}$ are the $t$ distinct $G$-conjugates of $\theta$; in particular, ${\mathbb{Q}}(\theta_{i})={\mathbb{Q}}(\theta)$ for all $i$. Now let $\Gamma:=\mathrm{Gal}(L/K)$, where $L:=\mathbb{F}(\chi,\theta)$ and $K:=\mathbb{F}(\chi)$, and let $\Gamma_{p}$ denote the Sylow $p$-subgroup of $\Gamma$. Then, any $\sigma\in\Gamma_{p}$ fixes $\chi_{N}$, and so it permutes $\theta_{1},...,\theta_{t}$. Thus the $p$-group $\Gamma_{p}$ acts on the set $\{\theta_{1},\ldots,\theta_{t}\}$ of $p^{\prime}$-size, and hence $\Gamma_{p}$ fixes some $\theta_{j}$. Setting $M:=\mathbb{F}(\theta_{j})=\mathbb{F}(\theta)$, we see that $MK$ is contained in the fixed field $L^{\Gamma_{p}}$ which has $p^{\prime}$-degree $\Gamma:\Gamma_{p}]$ over $K$ by Galois correspondence. It follows that $p\nmid[MK:K]=[M:K\cap M]=[\mathbb{F}(\theta):(\mathbb{F}(\chi)\cap\mathbb{F}(\theta))]$.
The same proof using Sylow subgroups of $\mathrm{Gal}(\mathbb{F}(\theta,\chi_{N})/\mathbb{F}(\chi_{N}))$ yields the fourth statement.
(ii) The assumption implies that all the characters $\gamma$ in $\operatorname{Irr}(J|\theta)$ have the same ramification index $e$: $\gamma_{N}=e\theta$, and so, by the Clifford correspondence [I76, (6.11)], if $\operatorname{Irr}(G|\theta)=\{\chi_{1},\ldots,\chi_{s}\}$ then
$\theta^{G},\chi_{i}]_{G}=[\theta,(\chi_{i})_{N}]_{N}=e$
for all $i$. Hence,
(3.2) $\theta^{G}=e\sum^{s}_{i=1}\chi_{i},$
and thus
$\bigl(\theta^{G}\bigr)_{N}=e\sum^{s}_{i=1}(\chi_{i})_{N}=e^{2}s\sum^{t}_{j=1}\theta_{j}$
in the notation of (i). It follows that $e^{2}st=|G/N|$, and so $p\nmid s$.
Now let $\Gamma:=\mathrm{Gal}(L/K)$, where $L:=\mathbb{F}(\chi_{1},\ldots,\chi_{s},\theta)$ and $K:=\mathbb{F}(\theta)$, and let $\Gamma_{p}$ denote the Sylow $p$-subgroup of $\Gamma$. Then, any $\sigma\in\Gamma_{p}$ fixes $\theta^{G}$, and so it permutes $\chi_{1},...,\chi_{s}$ by (3.2). Thus the $p$-group $\Gamma_{p}$ acts on the set $\{\chi_{1},\ldots,\chi_{s}\}$ of $p^{\prime}$-size, and hence $\Gamma_{p}$ fixes some $\chi=\chi_{i}$. Setting $M:=\mathbb{F}(\chi)$, we see that $MK$ is contained in the fixed field $L^{\Gamma_{p}}$ which has $p^{\prime}$-degree $\Gamma:\Gamma_{p}]$ over $K$ by Galois correspondence. It follows that $p\nmid[MK:K]=[M:K\cap M]=[\mathbb{F}(\chi):(\mathbb{F}(\chi)\cap\mathbb{F}(\theta))]$. ∎
Now we can prove the following “going-up and going-down” result:
Theorem 3.8.
Let $p$ be a fixed prime, $G$ a finite group, and $N$ a normal subgroup of $G$ such that $p\nmid|G/N|$. Then the following statements hold.
  1. (i)
    Suppose Conjecture B holds for all irreducible characters of NN at the prime $p$. Then Conjecture B holds for all irreducible characters of $G$ at the prime $p$.
  2. (ii)
    Suppose that $G/N$ is solvable and that Conjecture B holds for all irreducible characters of $G$ at the prime $p$. Then Conjecture B holds for all irreducible characters of NN at the prime $p$.
Proof.
(i) It suffices to prove the theorem in the case $\chi\in\operatorname{Irr}(G)$ has $a:=\nu_{p}(c(\chi))\geq 2$. By Proposition 3.7(i), for any $\theta\in\operatorname{Irr}(N)$ lying under $\chi$ we have
$\nu_{p}(c(\theta))=a,$
furthermore,
$p\nmid[M:K]$
with $M:={\mathbb{Q}}_{p}(\theta)$ and $K:={\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta)$. Since
$=[{\mathbb{Q}}_{p^{a}}M:{\mathbb{Q}}_{p^{a}}]/[{\mathbb{Q}}_{p^{a}}K:{\mathbb{Q}}_{p^{a}}]$
$=[M:({\mathbb{Q}}_{p^{a}}\cap M)]/[K:({\mathbb{Q}}_{p^{a}}\cap K)]$
$=[M:K]/[({\mathbb{Q}}_{p^{a}}\cap M):({\mathbb{Q}}_{p^{a}}\cap K)]$
is an integer, it follows that
$p\nmid[({\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\theta)):({\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta))].$
But ${\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta)\supseteq{\mathbb{Q}}_{p}$ and ${\mathbb{Q}}_{p^{a}}:{\mathbb{Q}}_{p}]=p^{a-1}$, so in fact we have shown that
${\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\theta)\subseteq{\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\chi).$
Denoting $p^{c}:=[{\mathbb{Q}}_{p^{a}}:({\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\theta))]$ and $p^{d}:=[{\mathbb{Q}}_{p^{a}}:({\mathbb{Q}}_{p^{a}}\cap{\mathbb{Q}}_{p}(\chi))]$, we now have
(3.3) $c\geq d.$
By hypothesis, $p^{c}|\theta(1)$. Next, $p\nmid|G/N|$ implies that $p\nmid\chi(1)/\chi(1)$, and so $p^{c}|\chi(1)$. Hence $p^{d}|\chi(1)$ by (3.3), and we are done.
(ii) It suffices to prove the theorem in the case $G/N$ is cyclic, and furthermore $\theta\in\operatorname{Irr}(N)$ has $b=\nu_{p}(c(\theta))\geq 2$. By Proposition 3.7(i), for any $\chi\in\operatorname{Irr}(G|\theta)$ we have
$\nu_{p}(c(\chi))=b.$
Since $\mathrm{Stab}_{G}(\theta)/N$ is cyclic in our case, using Proposition 3.7(ii) with $\mathbb{F}:={\mathbb{Q}}_{p}$, there exists some $\chi\in\operatorname{Irr}(G|\theta)$ such that
$p\nmid[{\mathbb{Q}}_{p}(\chi):({\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta))].$
Arguing as in (i), we see that
$p\nmid[({\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\chi)):({\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\chi)\cap{\mathbb{Q}}_{p}(\theta))],$
and hence
${\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\chi)\subseteq{\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\theta).$
Denoting $p^{c}:=[{\mathbb{Q}}_{p^{b}}:({\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\chi))]$ and $p^{d}:=[{\mathbb{Q}}_{p^{b}}:({\mathbb{Q}}_{p^{b}}\cap{\mathbb{Q}}_{p}(\theta))]$, we now have
(3.4) $c\geq d.$
By hypothesis, $p^{c}|\chi(1)$. Next, $p\nmid|G/N|$ implies that $p\nmid\chi(1)/\theta(1)$, and so $p^{c}|\theta(1)$. Hence $p^{d}|\theta(1)$ by (3.4), and we are done. ∎
Theorem 3.9.
The following statements hold.
  1. (i)
    Suppose that $S=\operatorname{SL}_{n}(q)$ and $p\nmid\gcd(n,q-1)$. Then Theorem D at the prime $p$ holds for $S=\operatorname{SL}_{n}(q)$ and for any finite group $L\rhd S$ with $p\nmid|L/S|$.
  2. (ii)
    Suppose that $S=\operatorname{SU}_{n}(q)$ and $p\nmid\gcd(n,q+1)$. Then Theorem D at the prime $p$ holds for $S=\operatorname{SU}_{n}(q)$ and for any finite group $L\rhd S$ with $p\nmid|L/S|$.
Proof.
(i) Embed $S$ in $G:=\operatorname{GL}_{n}(q)$, and consider $H:={\mathbf{Z}}(G)S$. Since $G/H\cong\mathsf{C}_{\gcd(n,q-1)}$, by Theorems 3.6 and 3.8(ii), Conjecture B holds for $H$ at $p$. Note that
$H={\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))S\times{\mathbf{O}}_{p}({\mathbf{Z}}(G))$
is a direct product. Indeed, suppose that $\alpha\cdot\mathrm{Id}=\beta g$ where $\alpha\cdot\mathrm{Id}\in{\mathbf{O}}_{p}({\mathbf{Z}}(G))$, $\beta\cdot\mathrm{Id}\in{\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))$, and $g\in S$. So $\alpha\in\mathbb{F}_{q}^{\times}$ is a $p$-element and $\beta\in\mathbb{F}_{q}^{\times}$ is a $p^{\prime}$-element. On the other hand,
$\alpha^{n}={\rm det}(\alpha\cdot\mathrm{Id})={\rm det}(\beta g)=\beta^{n},$
so $\alpha^{n}$ is both a $p$-element and a $p^{\prime}$-element, whence $\alpha^{n}=1$. But $\alpha^{q-1}=1$, so the order of the $p$-element $\alpha$ divides $\gcd(n,q-1)$. As $p\nmid\gcd(n,q-1)$, we conclude $\alpha=1$.
As ${\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))S$ is a direct factor of $H$, Conjecture B holds for ${\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))S$. Since $S$ is a normal subgroup of ${\mathbf{O}}_{p^{\prime}}({\mathbf{Z}}(G))S$ with quotient a cyclic $p^{\prime}$-group, so Conjecture B holds for $S$ at $p$ by Theorem 3.8(ii). By Theorem 3.8(i), Conjecture B then holds for $L$ at $p$.
(ii) The arguments are entirely similar to the ones in (i). ∎
Proof of Theorem D.
It follows from Theorems 3.2, 3.3, 3.6, and 3.9. ∎
4. Theorem E
For the definition of the inductive Feit condition, we refer the reader to [BKNT25, §3].
Theorem 4.1.
Let $G$ be a finite group and $\chi\in\operatorname{Irr}(G)$ be non-linear. Assume that simple groups involved in $G$ satisfy the inductive Feit condition. Then there exist $H<G$ and $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$ and ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$ divides $|G:H|$.
Proof.
Suppose that $N\trianglelefteq\,G$, $\theta\in{\rm Irr}(N)$ is an irreducible constituent of $\chi_{N}$ and let $\eta\in{\rm Irr}(T|\theta)$ be the Clifford correspondent of $\chi$ over $\theta$. By Lemma 2.1, we have that ${\mathbb{Q}}(\eta):{\mathbb{Q}}(\chi)]$ divides $|G:T|$. If $T<G$, then we are done. Hence, we may assume that $\chi$ is quasi-primitive. By [BKNT25, Theorem 4.7], there exists $H<G$ such that ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\psi)$, and we are done. (Notice that this theorem is stated for primitive characters, but as written after, it also holds for quasi-primitive characters.) ∎
The following is Theorem E.
Corollary 4.2.
Suppose that the finite simple groups satisfy the inductive Feit condition. Then the following hold:
  1. (a)
    If $G$ is a finite group and $\chi\in{\rm Irr}(G)$, then $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$.
  2. (b)
    Let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension, let $p$ be a prime, and $b\in{\mathbb{Z}}_{\geq 0}$. Then $\mathbb{F}={\mathbb{Q}}(\chi)$ for some $\chi\in{\rm Irr}(G)$, where $G$ is a finite group with $\nu_{p}(|G|)=b$, if and only if
    $\nu_{p}(c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b.$
Proof.
We argue by induction on $|G|$. If $\chi$ is linear, then $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]=c(\chi)$ divides the exponent of $G$, and we are done. We may assume that $\chi$ is not linear. By Theorem 4.1, there exists $H<G$ and $\psi\in\operatorname{Irr}(H)$ such that ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$ and ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$ divides $|G:H|$. By induction, we have that $c(\psi)[{\mathbb{Q}}_{c(\psi)}:{\mathbb{Q}}(\psi)]$ divides $|H|$. Since ${\mathbb{Q}}(\chi)\subseteq{\mathbb{Q}}(\psi)$, we have that $c(\chi)$ divides $c(\psi)$. Let $K={\mathbb{Q}}_{c(\chi)}\cap{\mathbb{Q}}(\psi)$. By elementary Galois theory, $c(\chi)[{\mathbb{Q}}_{c(\chi)}:K]$ divides $c(\psi)[{\mathbb{Q}}_{c(\psi)}:{\mathbb{Q}}(\psi)]$, which divides $|H|$. Also, $K:{\mathbb{Q}}(\chi)]$ divides ${\mathbb{Q}}(\psi):{\mathbb{Q}}(\chi)]$, which divides $|G:H|$, and part (a) follows.
Now, let $\mathbb{F}/{\mathbb{Q}}$ be an abelian extension, let $p$ be a prime, and $b\in{\mathbb{Z}}_{\geq 0}$ such that $\nu_{p}(c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}])\leq b$. By Lemma 2.3, there exists a group $M$ of order $c(\mathbb{F})[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]$ and $\varphi\in\operatorname{Irr}(M)$ such that ${\mathbb{Q}}(\varphi)=F$. The direct product $G:=M\times C$, where $C$ is the cyclic $p$-group of order $p^{b}/(c(\mathbb{F})_{p}[{\mathbb{Q}}_{c(\mathbb{F})}:\mathbb{F}]]_{p})$, and the character $\chi:=\varphi\otimes\textbf{1}_{C}\in\operatorname{Irr}(G)$ satisfy the required condition. Conversely, assume that $G$ is a finite group such that $|G|_{p}=p^{b}$ and $\chi\in{\rm Irr}(G)$. By part (a), we have that $c(\chi)[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $|G|$, which implies that $c(\chi)_{p}[{\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]_{p}\leq|G|_{p}=p^{b}$, and we are done. ∎
For solvable groups, it was proven in [C88, Theorem 0.1] that in fact ${\mathbb{Q}}_{c(\chi)}:{\mathbb{Q}}(\chi)]$ divides $\chi(1)$, but this is not true in general (as shown, for instance, by ${\mathrm{Alt}}_{5}$ and the irreducible characters of degree 3).